### v2

# Nonsingular black holes and spherically symmetric objects in nonlinear electrodynamics with a scalar field

Abstract
In general relativity with vector and scalar fields given by the Lagrangian ${\cal L}(F,\phi,X)$, where $F$ is a Maxwell term and $X$ is a kinetic term of the scalar field $\phi$, we study the linear stability of static and spherically symmetric objects without curvature singularities at their centers. We show that the background solutions are generally described by either purely electrically or magnetically charged objects with a nontrivial scalar-field profile. In theories with the Lagrangian ~$\tilde{{\cal L}}(F)+K(\phi,X)$, which correspond to nonlinear electrodynamics with a k-essence scalar field, angular Laplacian instabilities induced by vector-field perturbations exclude all the regular spherically symmetric solutions including nonsingular black holes. In theories described by the Lagrangian ${\cal L}=X+\mu(\phi)F^{n}$, where $\mu$ is a function of $\phi$ and $n$ is a constant, the absence of angular Laplacian instabilities of spherically symmetric objects requires that $n>1/2$, under which nonsingular black holes with apparent horizons are not present. However, for some particular ranges of $n$, there are horizonless compact objects with neither ghosts nor Laplacian instabilities in the small-scale limit. In theories given by ${\cal L}=X\kappa(F)$, where $\kappa$ is a function of $F$, regular spherically symmetric objects are prone to Laplacian instabilities either around the center or at spatial infinity. Thus, in our theoretical framework, we do not find any example of linearly stable nonsingular black holes.
preprint: YITP-24-167, WUCG-24-10
I Introduction
General Relativity (GR) is a fundamental pillar for describing gravitational interactions in both strong and weak field regimes. The vacuum solution to the Einstein equation on a static and spherically symmetric (SSS) background is described by a Schwarzschild line metric that contains a mass $M$ of the source. In Einstein-Maxwell theory with the electromagnetic Lagrangian $F=-F_{\mu\nu}F^{\mu\nu}/4$, where $F_{\mu\nu}=\partial_{\mu}A_{\nu}-\partial_{\nu}A_{\mu}$ is the Maxwell tensor with a vector field $A_{\mu}$, the resulting solution is given by a Reissner-Nordström (RN) metric with electric or magnetic charges. For both Schwarzschild and RN black holes (BHs), there are singularities at the origin ($r=0$) with divergent curvature quantities. This divergent property at $r=0$ also persists for rotating BHs present in the framework of GR.
In GR, Penrose’s singularity theorem [1] establishes that BH singularities at the origin can arise as a natural consequence of gravitational collapse. The validity of this theorem, however, hinges on several assumptions regarding the structure of spacetime and the properties of matter. Among these is the requirement of global hyperbolicity of spacetime. Violating this condition can potentially lead to the existence of nonsingular BHs.
A notable example of such a solution was first introduced by Bardeen [2], who proposed a nonsingular BH with metric components that remain finite as $r\to 0$. Since then, various other regular BH metrics have been proposed in the literature, which offers alternative frameworks for addressing the singularity problem in BH physics [3, 4, 5, 6, 7, 8, 9].
Even though the nonsingular metrics are given apriori in the aforementioned approach, it remains to be seen whether they can be realized in some concrete theories. For this purpose, we need to take into account additional degrees of freedom (DOFs) beyond those appearing in GR. For example, we may consider scalar-tensor theories in which a new scalar DOF is incorporated into the gravitational action [10]. In most general scalar-tensor theories with second-order field equations of motion (Horndeski theories [11, 12, 13, 14]), it is known that the existence of SSS asymptotically-flat hairy BH solutions without ghost/Laplacian instabilities is quite limited [15, 16, 17, 18]. For a radial dependent scalar profile $\phi(r)$, we need a coupling between $\phi$ and a Gauss-Bonnet curvature invariant [19, 20, 21, 22, 23, 24], but the Gauss-Bonnet term diverges at $r=0$. Hence the construction of nonsingular BHs in the context of scalar-tensor theories is generally challenging.
If we consider vector-tensor theories in nonlinear electrodynamics (NED) given by the Lagrangian ${\cal L}(F)$, where ${\cal L}$ is a nonlinear function of $F$, it is possible to realize nonsingular BHs without curvature singularities at $r=0$ [25, 26, 27, 28, 4, 7, 29, 30]. The NED Lagrangian accommodates Euler-Heisenberg theory [31] as well as Born-Infeld theory [32]. In such subclasses of NED theories, the resulting SSS BH solutions possess curvature singularities at $r=0$ [33, 34, 35, 36]. However, there are nonsingular electrically or magnetically charged BHs for some specific choices of the NED Lagrangian. Thus, the vector field with nonlinear Lagrangians of $F$ allows an interesting possibility for realizing regular BHs even at the classical level.
To determine the stability of nonsingular BHs, it is essential to analyze their linear stability by using BH perturbation theory. In Refs. [37, 38, 39, 40, 41], the authors discussed the BH stability by considering the propagation of dynamical perturbations in the region outside the event horizon. Although the conditions for the absence of ghosts and Laplacian instabilities can be satisfied outside the horizon, a recent analysis [42] shows that there is angular Laplacian instability of vector-field perturbations around the regular center. This instability manifests for both electric and magnetic BHs, leading to the rapid enhancement of metric perturbations. Consequently, the nonsingular background metric cannot be maintained as a steady-state solution. This means that nonsingular SSS BHs cannot be realized in the context of NED with the Lagrangian ${\cal L}(F)$.
Motivated by the no-go result in the context of NED, we extend our analysis to explore whether similar properties persist in more general classical field theories. To this end, we incorporate a scalar field $\phi$ with a kinetic term $X$ into the NED framework, considering a Lagrangian of the form ${\cal L}(F,\phi,X)$. For the gravity sector, we consider GR described by the Lagrangian $M_{\rm Pl}^{2}R/2$, where $M_{\rm Pl}$ is the reduced Planck mass and $R$ is the Ricci scalar. Applying such theories to the SSS background, we will show that there are no solutions with mixed electric and magnetic charges (as it happens in NED). Hence we can focus on either electrically or magnetically charged objects, with a nontrivial scalar-field profile.
The theories we will study in this paper belong to a subclass of scalar-vector-tensor theories with second-order field equations of motion. Since they respect $U(1)$ gauge symmetry, there are one scalar, two transverse vectors, and two tensor polarizations as the propagating DOFs. To derive the stability conditions of those five DOFs, we consider linear perturbations in both odd- and even-parity sectors on the SSS background. We expand the corresponding action up to the second order in perturbations by taking into account both electric and magnetic charges. For the electric case, a similar analysis was performed in Refs. [43, 44] as a subclass of Maxwell-Horndeski theories. Since the linear stability of magnetic SSS objects has not been addressed yet, we will do so in this paper.
After deriving conditions for the absence of ghosts and Laplacian instabilities of the five dynamical DOFs, we will apply them to three subclasses of ${\cal L}(F,\phi,X)$ theories: (i) ~$\tilde{\cal L}(F)+K(\phi,X)$, (ii) ${\cal L}=X+\mu(\phi)F^{n}$, and (iii) ${\cal L}=X\kappa(F)$. We will show that regular SSS objects realized by theories (i) and (iii), which include nonsingular BHs, are excluded by Laplacian instabilities of vector-field perturbations around the origin. In theories (ii), the absence of angular Laplacian instabilities requires the condition $n>1/2$, under which there are no regular BHs with apparent horizons. Thus, even by extending NED to more general theories with the Lagrangian ${\cal L}(F,\phi,X)$, we do not find even a single example of nonsingular BHs without instabilities.
This shows the general difficulty of constructing BHs without singularities in classical field theories. In theories (ii), however, we will show the existence of linearly stable SSS compact objects without apparent horizons. These regular solutions are present for both electric and magnetic configurations. We will clarify the regions of $n$ in which the regular horizonless compact objects are subject to neither ghosts nor Laplacian instabilities.
This paper is organized as follows. In Sec. II, we derive the SSS background solutions and discuss the properties of them for the electrically and magnetically charged cases. In Sec. III, we expand the action up to quadratic order in perturbations and obtain conditions under which neither ghosts nor Laplacian instabilities are present for five dynamical DOFs. In Sec. IV, we apply the linear stability conditions to theories (i) mentioned above and show that nonsingular SSS objects are prone to angular Laplacian instability. In Sec. V, we show that the absence of angular Laplacian instabilities demands the condition $n>1/2$, under which nonsingular BHs with apparent horizons do not exist. We also clarify the parameter space of $n$ in which horizonless regular compact objects suffer from neither ghosts nor Laplacian instabilities. In Sec. VI, we show that nonsingular SSS objects in theories (iii) are excluded by Laplacian instabilities either at the origin or at spatial infinity. Sec. VII is devoted to conclusions.
II Field equations on the SSS background
We consider theories in which the Lagrangian ${\cal L}$ in the matter sector depends on a scalar field $\phi$ and the two scalar products
$X\equiv-\frac{1}{2}\partial_{\mu}\phi\partial^{\mu}\phi\,,\qquad F\equiv-\frac{1}{4}F_{\mu\nu}F^{\mu\nu}\,,$ (1)
where $F_{\mu\nu}=\partial_{\mu}A_{\nu}-\partial_{\nu}A_{\mu}$ is the field strength of a covector field $A_{\mu}$. For the gravity sector, we consider GR described by the Einstein-Hilbert Lagrangian $M_{\rm Pl}^{2}R/2$. Then, the total action is given by
${\cal S}=\int{\rm d}^{4}x\sqrt{-g}\left[\frac{M_{\rm Pl}^{2}}{2}R+{\cal L}(F,\phi,X)\right]\,,$ (2)
where $g$ is a determinant of the metric tensor $g_{\mu\nu}$.
We study SSS solutions on the background given by the line element
${\rm d}s^{2}=-f(r){\rm d}t^{2}+h^{-1}(r){\rm d}r^{2}+r^{2}({\rm d}\theta^{2}+\sin^{2}\theta\,{\rm d}\varphi^{2})\,,$ (3)
where $f$ and $h$ are functions of the radial distance $r$. For the positivity of $-g=(f/h)r^{4}\sin^{2}\theta$ in Eq. (2), we require that $f/h$ is positive. For the later convenience, we introduce the $r$-dependent function $N(r)$ satisfying
$f(r)=N(r)h(r)\,,$ (4)
so that
$N(r)>0\,.$ (5)
In the following, we will use the two functions $N(r)$ and $h(r)$ instead of $f(r)$ and $h(r)$.
On the SSS background (3), we consider a radial dependent scalar-field profile $\phi(r)$. For the covector field $A_{\mu}$, the presence of $U(1)$ gauge symmetry in theories given by the action (2) allows us to express the vector-field components in the form
$A_{\mu}{\rm d}x^{\mu}=A_{0}(r){\rm d}t-q_{M}\cos\theta{\rm d}\varphi\,,$ (6)
where $A_{0}$ is a function of $r$, and $q_{M}$ is a constant corresponding to a magnetic charge. The scalar products defined in Eq. (1) reduce to
$X=-\frac{1}{2}h\phi^{\prime 2}\,,\qquad F=\frac{A_{0}^{\prime 2}}{2N}-\frac{q_{M}^{2}}{2r^{4}}\,,$ (7)
where a prime represents the derivative with respect to $r$.
Varying the action (2) with respect to $N$, $h$, $A_{0}$, and $\phi$, we obtain
$h^{\prime}-\frac{1-h}{r}-\frac{r}{M_{\rm Pl}^{2}N}\left(N{\cal L}-A_{0}^{\prime 2}{\cal L}_{,F}\right)=0\,,$ (8)
$\frac{N^{\prime}}{N}=\frac{r\phi^{\prime 2}{\cal L}_{,X}}{M_{\rm Pl}^{2}}\,,$ (9)
$\left(\frac{r^{2}A_{0}^{\prime}{\cal L}_{,F}}{\sqrt{N}}\right)^{\prime}=0\,,$ (10)
${\cal E}_{\phi}\equiv\left(\sqrt{N}hr^{2}\phi^{\prime}{\cal L}_{,X}\right)^{\prime}+\sqrt{N}r^{2}{\cal L}_{,\phi}=0\,,$ (11)
where the notations like ${\cal L}_{,F}\equiv\partial{\cal L}/\partial F$ are used for partial derivatives. If ${\cal L}_{,X}=0$, we require that $N^{\prime}=0$ in general, or $N(r)=1$ by fixing boundary conditions at spatial infinity. If $A_{0}^{\prime}\neq 0$, we can integrate Eq. (10) to give
${\cal L}_{,F}=\frac{q_{E}\sqrt{N}}{r^{2}A_{0}^{\prime}}\,,$ (12)
where $q_{E}$ is an integration constant corresponding to an electric charge. From Eqs. (8) and (12), we have
${\cal L}=\frac{M_{\rm Pl}^{2}}{r^{2}}\left(rh^{\prime}+h-1\right)+\frac{q_{E}A_{0}^{\prime}}{\sqrt{N}r^{2}}\,.$ (13)
Using Eqs. (9) and (11), we can express ${\cal L}_{,X}$ and ${\cal L}_{,\phi}$, as
${\cal L}_{,X}$ $=$ $\frac{M_{\rm Pl}^{2}N^{\prime}}{r\phi^{\prime 2}N}\,,$ (14)
${\cal L}_{,\phi}$ $=$ $-\frac{1}{\sqrt{N}r^{2}}\left(\frac{M_{\rm Pl}^{2}N^{\prime}hr}{\sqrt{N}\phi^{\prime}}\right)^{\prime}\,,$ (15)
which are valid for $\phi^{\prime}\neq 0$. Taking the $r$ derivative of Eq. (13), i.e., ${\cal L}^{\prime}(r)={\cal L}_{,F}F^{\prime}+{\cal L}_{,\phi}\phi^{\prime}+{\cal L}_{,X}X^{\prime}$, and employing Eqs. (7), (12), (14), and (15), we obtain
$4q_{E}r^{2}A_{0}^{\prime 2}-M_{\rm Pl}^{2}r^{2}N^{-3/2}[2N^{2}(r^{2}h^{\prime\prime}-2h+2)$
$+rN(2rhN^{\prime\prime}+3rh^{\prime}N^{\prime}+2hN^{\prime})-r^{2}hN^{\prime 2}]A_{0}^{\prime}$
$+4Nq_{E}q_{M}^{2}/r^{2}=0\,,$ (16)
so that $A_{0}^{\prime}$ is known algebraically in terms of $h$, $N$, and its $r$ derivatives. Interestingly, the $\phi$-dependent terms completely vanish in Eq. (16).
We will consider SSS objects that are regular at $r=0$, including nonsingular BHs with apparent horizons [2, 3, 4, 5, 6, 7, 8, 9]. In such cases, the Ricci scalar $R$, the squared Ricci tensor $R_{\mu\nu}R^{\mu\nu}$, and the squared Riemann tensor $R_{\mu\nu\rho\sigma}R^{\mu\nu\rho\sigma}$ do not diverge at $r=0$. This requires that $h$ and $N$ are expanded around $r=0$, as [45]
$h(r)$ $=$ $1+\sum_{n=2}^{\infty}h_{n}r^{n}\,,$ (17)
$N(r)$ $=$ $N_{0}+\sum_{n=2}^{\infty}N_{n}r^{n}\,,$ (18)
where $h_{n}$, $N_{0}$, and $N_{n}$ are constants. Note that $N_{0}$ is positive due to the condition (5). We substitute Eqs. (17)-(18) and their $r$ derivatives into Eq. (16). Solving the resulting equation for $A_{0}^{\prime}$ and expanding it around $r=0$, we find
$A_{0}^{\prime}=\pm\frac{\sqrt{N_{0}}\sqrt{-(q_{E}q_{M})^{2}}}{q_{E}r^{2}}+{\cal O}(r^{0})\,.$ (19)
Then, we have real solutions to $A_{0}^{\prime}$ only if
$q_{E}q_{M}=0\,,$ (20)
and hence the dyon BHs with mixed electric and magnetic charges are not allowed [^1] Equivalently, we may study the lowest order of a discriminant of the second-order algebraic equation and show that it is always negative.. In the following, we will separate the discussion into the electrically and magnetically charged cases.
II.1 Electric case
For $q_{E}\neq 0$ and $q_{M}=0$, the nonvanishing solution to Eq. (16) is given by
$A_{0}^{\prime}$ $=$ $M_{\rm Pl}^{2}[2N^{2}(r^{2}h^{\prime\prime}-2h+2)+rN(2rhN^{\prime\prime}+3rh^{\prime}N^{\prime}$ (21)
$+2hN^{\prime})-r^{2}hN^{\prime 2}]/(4N^{3/2}q_{E})\,,$
which depends on the background metrics $h$ and $N$. Using the expansions (17) and (18) around $r=0$, we have
$A_{0}^{\prime}=\frac{2M_{\rm Pl}^{2}N_{2}}{\sqrt{N_{0}}q_{E}}r^{2}+\frac{M_{\rm Pl}^{2}(4N_{0}h_{3}+9N_{3})}{2\sqrt{N_{0}}q_{E}}r^{3}+{\cal O}(r^{4})\,,$ (22)
which is finite at $r=0$. Substituting Eq. (21) into Eqs. (7) and (13), we know $F$ and ${\cal L}$ in terms of $h$, $N$, and their $r$ derivatives. Under the expansions (17) and (18) around $r=0$, it follows that
$F$ $=$ $\frac{2M_{\rm Pl}^{4}N_{2}^{2}}{N_{0}^{2}q_{E}^{2}}r^{4}+\frac{M_{\rm Pl}^{4}N_{2}(4N_{0}h_{3}+9N_{3})}{N_{0}^{2}q_{E}^{2}}r^{5}$ (23)
$+{\cal O}(r^{6})\,,$
${\cal L}$ $=$ $\frac{M_{\rm Pl}^{2}(3N_{0}h_{2}+2N_{2})}{N_{0}}+\frac{3M_{\rm Pl}^{2}(4N_{0}h_{3}+3N_{3})}{2N_{0}}r$ (24)
$+{\cal O}(r^{2})\,,$
which are both finite at $r=0$.
II.2 Magnetic case
For $q_{M}\neq 0$ and $q_{E}=0$, the solution to Eq. (16) yields
$A_{0}^{\prime}=0\,.$ (25)
From Eq. (13), we have
${\cal L}=\frac{M_{\rm Pl}^{2}}{r^{2}}\left(rh^{\prime}+h-1\right)\,.$ (26)
Applying the expansion (17) to Eq. (26) gives
${\cal L}=3M_{\rm Pl}^{2}h_{2}+4M_{\rm Pl}^{2}h_{3}r+{\cal O}(r^{2})\,,$ (27)
which approaches a constant as $r\to 0$. We note that the quantity $F=-q_{M}^{2}/(2r^{4})$ diverges at $r=0$, but the form of ${\cal L}(F,\phi,X)$ can be designed to have the regular behavior (27) at the origin. Indeed, this was already shown for nonsingular magnetic BHs present in the pure NED described by the Lagrangian ${\cal L}(F)$ [27, 28].
III Perturbations on the SSS background
The linear stability of BHs can be analyzed by considering perturbations on the SSS background (3) [46, 47, 48, 49]. We write the metric tensor in the form $g_{\mu\nu}=\bar{g}_{\mu\nu}+h_{\mu\nu}$, where $\bar{g}_{\mu\nu}$ is the background value and $h_{\mu\nu}$ is the metric perturbation. We expand $h_{\mu\nu}$ in terms of the spherical harmonics $Y_{lm}(\theta,\varphi)$. Without loss of generality, we will focus on the mode $m=0$ and express $Y_{l0}$ as $Y_{l}$ in the following. We also omit the summation for $l$ for each perturbed variable.
We choose the four gauge conditions $h_{t\theta}=0$, $h_{\theta\theta}=0$, $h_{\varphi\varphi}=0$, and $h_{\theta\varphi}=0$. In this case, the four components of $\xi^{\mu}$ under the infinitesimal coordinate transformation $x^{\mu}\to x^{\mu}+\xi^{\mu}$ are fixed. Then, the components of $h_{\mu\nu}$ are given by [50, 51, 52, 44]
$h_{tt}=f(r)H_{0}(t,r)Y_{l}(\theta),\quad h_{tr}=H_{1}(t,r)Y_{l}(\theta),\quad h_{t\theta}=0,$
$h_{t\varphi}=-Q(t,r)(\sin\theta)Y_{l,\theta}(\theta),\;\;h_{rr}=f^{-1}(r)H_{2}(t,r)Y_{l}(\theta),$
$h_{r\theta}=h_{1}(t,r)Y_{l,\theta}(\theta),\quad h_{r\varphi}=-W(t,r)(\sin\theta)Y_{l,\theta}(\theta),$
$h_{\theta\theta}=0,\quad h_{\varphi\varphi}=0,\quad h_{\theta\varphi}=0\,,$ (28)
where $H_{0}$, $H_{1}$, $H_{2}$, $h_{1}$, $Q$, and $W$ are functions of $t$ and $r$.
We also decompose the scalar and vector fields, as
$\phi$ $=$ $\bar{\phi}(r)+\delta\phi(t,r)Y_{l}(\theta),$ (29)
$A_{\mu}$ $=$ $\bar{A}_{\mu}(r)+\delta A_{\mu}\,,$ (30)
where
$\delta A_{t}=\delta A_{0}(t,r)Y_{l}(\theta),\qquad\delta A_{r}=\delta A_{1}(t,r)Y_{l}(\theta),\qquad$
$\delta A_{\theta}=0,\qquad\delta A_{\varphi}=-\delta A(t,r)(\sin\theta)Y_{l,\theta}(\theta)\,.$ (31)
Here, we have set $\delta A_{\theta}=0$ by exploiting the fact that the action (2) respects $U(1)$ gauge invariance.
The odd-parity sector contains three perturbations $Q$, $W$, and $\delta A$, while there are seven perturbed fields $H_{0}$, $H_{1}$, $H_{2}$, $h_{1}$, $\delta\phi$, $\delta A_{0}$, and $\delta A_{1}$ in the even-parity sector. After integrating out all nondynamical fields, we have one scalar perturbation $\delta\phi$, two vector modes arising from $\delta A_{\mu}$, and two tensor polarizations arising from the gravity sector. For the electric BH, the linear stability conditions of such five dynamical perturbations were derived in Refs. [43, 44] for more general Maxwell-Horndeski theories. For the magnetic BH, the stability issue has not been addressed yet in theories given by the action (2). In the following, we will obtain the full second-order action of linear perturbations and study the stability of SSS objects with electric and magnetic charges, in turn.
III.1 Second-order action
We expand the action (2) up to quadratic order in perturbations and integrate it with respect to $\theta$ and $\varphi$. After the integration by parts, the second-order action can be expressed in the form
${\cal S}^{(2)}=\int{\rm d}t{\rm d}r\,\left({\cal L}_{1}+{\cal L}_{2}\right)\,,$ (32)
where
${\cal L}_{1}$ $=$ $a_{0}H_{0}^{2}+H_{0}\left[a_{1}H_{2}^{\prime}+La_{2}h_{1}^{\prime}+(a_{3}+La_{4})H_{2}+La_{5}h_{1}+La_{6}\delta A+a_{7}\delta\phi^{\prime}+a_{8}\delta\phi\right]+Lb_{1}H_{1}^{2}$ (33)
$+H_{1}(b_{2}\dot{H}_{2}+Lb_{3}\dot{h}_{1}+b_{4}\dot{\delta\phi})+c_{0}H_{2}^{2}+LH_{2}(c_{1}h_{1}+c_{2}\delta A)+c_{3}H_{2}\delta\phi^{\prime}+c_{4}H_{2}\delta\phi+L(d_{0}\dot{h}_{1}^{2}+d_{1}h_{1}^{2})$
$+Lh_{1}(d_{2}\delta A_{0}+d_{3}\delta A^{\prime}+d_{4}\delta\phi)+s_{1}(\delta A_{0}^{\prime}-\dot{\delta A_{1}})^{2}+(s_{2}H_{0}+s_{3}H_{2}+Ls_{4}\delta A+s_{5}\delta\phi^{\prime}+s_{6}\delta\phi)(\delta A_{0}^{\prime}-\dot{\delta A_{1}})$
$+L(s_{7}\delta A_{0}^{2}+s_{8}\delta A_{1}^{2})+u_{1}\dot{\delta\phi}^{2}+u_{2}\delta\phi^{\prime 2}+\left(Lu_{3}+\tilde{u}_{3}\right)\delta\phi^{2}+u_{4}\delta\phi^{\prime}\delta A+u_{5}\delta\phi\,\delta A\,,$
${\cal L}_{2}$ $=$ $L[p_{1}(r\dot{W}-rQ^{\prime}+2Q)^{2}+p_{2}\delta A(r\dot{W}-rQ^{\prime}+2Q)+p_{3}\dot{\delta A^{2}}+p_{4}\delta A^{\prime 2}+Lp_{5}\delta A^{2}+(Lp_{6}+p_{7})W^{2}$ (34)
$~{}~{}+(Lp_{8}+p_{9})Q^{2}+p_{10}Q\delta A_{0}+p_{11}Qh_{1}+p_{12}W\delta A_{1}]\,,$
where a dot represents the derivative with respect to $t$, the coefficients $a_{0}$ etc are given in Appendix A, and
$L\equiv l(l+1)\,.$ (35)
For both electric and magnetic BHs, we have
$s_{4}=-\frac{q_{M}A_{0}^{\prime}{\cal L}_{,FF}}{\sqrt{N}\,r^{2}}=0\,,$ (36)
whose condition will be used in the following.
In the odd-parity sector, there are two dynamical perturbations [44, 42]
$\chi\equiv r\dot{W}-rQ^{\prime}+2Q-\frac{2{\cal L}_{,F}rA_{0}^{\prime}}{M_{\rm Pl}^{2}}\delta A\,,$ (37)
$\delta A\,,$ (38)
which correspond to the gravitational and vector-field perturbations, respectively.
In the even-parity sector, there are three dynamical fields given by
$\psi\equiv rH_{2}-Lh_{1}\,,$ (39)
$V\equiv\delta A_{0}^{\prime}-\dot{\delta A}_{1}+\frac{s_{2}H_{0}+s_{3}H_{2}+s_{5}\delta\phi^{\prime}+s_{6}\delta\phi}{2s_{1}},$ (40)
$\delta\phi\,,$ (41)
which correspond to the gravitational, vector-field, and scalar-field perturbations, respectively.
Although we have split perturbations into even- and odd-parity modes, the way they couple with each other depends on the background. For instance, if $q_{M}$ does not vanish, there are coupling terms between $\delta\phi$ and $\delta A$, as $u_{4}\neq 0\neq u_{5}$.
To simplify the analysis of the propagating dynamical DOFs, we introduce two auxiliary fields, $V$ and $\chi$, allowing the action to be reformulated as follows:
$\tilde{\cal S}^{(2)}=\int{\rm d}t{\rm d}r\,\left(\tilde{{\cal L}}_{1}+\tilde{{\cal L}}_{2}\right)\,,$ (42)
where
$\tilde{{\cal L}_{1}}$ $=$ ${\cal L}_{1}-s_{1}\biggl{[}\delta A_{0}^{\prime}-\dot{\delta A}_{1}$ (43)
$\qquad\quad+\frac{s_{2}H_{0}+s_{3}H_{2}+s_{5}\delta\phi^{\prime}+s_{6}\delta\phi}{2s_{1}}-V\biggr{]}^{2},$
~$\tilde{{\cal L}}_{2}$ $=$ ${\cal L}_{2}-Lp_{1}\left[r\dot{W}-rQ^{\prime}+2Q-\frac{2{\cal L}_{,F}rA_{0}^{\prime}}{M_{\rm Pl}^{2}}\delta A-\chi\right]^{2}.$
Varying the action (42) with respect to $V$ and $\chi$, we obtain the relations (40) and (37), respectively. Then, we find that the action (42) is equivalent to (32).
III.2 Stability of electric SSS objects
We first study the linear stability of electric SSS objects ($q_{M}=0$) by exploiting the action (42). This issue was addressed in Refs. [43, 44] as a special case of Maxwell-Horndeski theories, but we will revisit it to make a comparison with the stability of magnetic SSS objects. For electric objects, ~$\tilde{{\cal L}}_{1}$ is composed of even-parity perturbations alone, while ~$\tilde{{\cal L}}_{2}$ contains only odd-parity perturbations. In the following, we will study the case $h>0$, but we will also address the case $h<0$ at the end of this subsection.
III.2.1 Odd-parity sector
We first consider perturbations in the odd-parity sector. Varying the Lagrangian ~$\tilde{{\cal L}}_{2}$ with respect to $Q$ and $W$, we obtain
$Q$ $=$ $-\frac{h[2rN\chi^{\prime}+(2N-rN^{\prime})\chi]}{2(L-2)N}\,,$ (45)
$W$ $=$ $-\frac{r\dot{\chi}}{(L-2)Nh}\,.$ (46)
Substituting these expressions of $Q$ and $W$ and their $t$, $r$ derivatives into ~$\tilde{{\cal L}}_{2}$ and integrating it by parts, the second-order Lagrangian is expressed in the form
~$\tilde{{\cal L}}_{2}=\dot{\vec{\Psi}}_{\rm A}^{t}{\bm{K}}_{\rm A}\dot{\vec{\Psi}}_{\rm A}+\vec{\Psi}^{\prime t}{\bm{G}}_{\rm A}\vec{\Psi}_{\rm A}^{\prime}+\vec{\Psi}^{t}{\bm{M}}_{\rm A}\vec{\Psi}_{\rm A}\,,$ (47)
where ${\bm{K}}_{\rm A}$, ${\bm{G}}_{\rm A}$, and ${\bm{M}}_{\rm A}$ are the $2\times 2$ matrices, and
$\vec{\Psi}_{\rm A}^{t}=\left(\chi,\delta A\right)\,.$ (48)
The matrix ${\bm{K}}_{\rm A}$ has only the diagonal components $(K_{\rm A})_{11}$ and $(K_{\rm A})_{22}$, so that the no-ghost conditions are given by
$(K_{\rm A})_{11}$ $=$ $\frac{M_{\rm Pl}^{2}L}{4hN^{3/2}(L-2)}>0\,,$ (49)
$(K_{\rm A})_{11}(K_{\rm A})_{22}$ $=$ $\frac{M_{\rm Pl}^{2}L^{2}{\cal L}_{,F}}{8h^{2}N^{2}(L-2)}>0\,.$ (50)
The inequality (49) automatically holds for $l\geq 2$, while the other inequality (50) is satisfied if
${\cal L}_{,F}>0\,.$ (51)
To study the propagation of dynamical perturbations along the radial direction, we first vary the Lagrangian (47) with respect to $\chi$ and $\delta A$. Then, we assume solutions to the perturbation equations in the WKB form
$\vec{\Psi}_{\rm A}^{t}=(\vec{\Psi}_{0}^{t})_{\rm A}e^{-i(\omega t-kr)}\,,$ (52)
where $(\vec{\Psi}_{0}^{t})_{\rm A}=(\chi_{0},\delta A_{0})$ is a constant vector, $\omega$ is an angular frequency, and $k$ is a wavenumber. This gives the algebraic equation ${\bm{U}}_{\rm A}(\vec{\Psi}_{0})_{\rm A}=0$, where ${\bm{U}}_{\rm A}$ is a $2\times 2$ matrix. To allow the existence of nonvanishing solutions to $(\vec{\Psi}_{0})_{\rm A}$, we require that the determinant of ${\bm{U}}_{\rm A}$ vanishes, i.e.,
${\rm det}~{}{\bm{U}}_{\rm A}=0\,.$ (53)
In the regime $h>0$, the radial propagation speed $c_{r}=h^{-1/2}{\rm d}r/{\rm d}\tau$ in proper time $\tau=\int\sqrt{Nh}\,{\rm d}t$ can be derived by substituting $\omega=h\sqrt{N}c_{r}k$ into Eq. (53). Taking the large $\omega$ and $k$ limits, we obtain the two solutions
$c_{r}^{2}=1\,,\quad{\rm for}\quad\vec{\Psi}_{\rm A}^{t}=\left(\chi,\delta A\right)\,,$ (54)
so that the two dynamical fields $\chi$ and $\delta A$ propagate with the speed of light along the radial direction.
The angular propagation speed measured by the proper time $\tau$ is given by $c_{\Omega}=r{\rm d}\theta/{\rm d}\tau=(r/\sqrt{Nh})(\omega/l)$. Taking the large $\omega$ and $l$ limits in Eq. (53), we obtain the following two solutions
$c_{\Omega}^{2}=1\,,\quad{\rm for}\quad\vec{\Psi}_{\rm A}^{t}=\left(\chi,\delta A\right)\,,$ (55)
and hence the angular propagation speeds of both $\chi$ and $\delta A$ are luminal.
III.2.2 Even-parity sector
In the even-parity sector, the product $H_{0}^{2}$ present in ${\cal L}_{1}$ disappears in ~$\tilde{\cal L}_{1}$ as a result of introducing the Lagrange multiplier $V$. Since ~$\tilde{\cal L}_{1}$ only contains terms linear in $H_{0}$, varying ~$\tilde{\cal L}_{1}$ with respect to $H_{0}$ puts constraints on other perturbed fields. We use this equation to express $h_{1}$ in terms of $\psi$, $V$, $\delta\phi$, and their derivatives. We also vary ~$\tilde{\cal L}_{1}$ with respect to $H_{1}$, $\delta A_{0}$, $\delta A_{1}$ and eliminate these fields from ~$\tilde{\cal L}_{1}$ by using their equations of motion. Up to boundary terms, we can express ~$\tilde{\cal L}_{1}$ in the following form
~$\tilde{{\cal L}}_{1}=\dot{\vec{\Psi}}_{\rm B}^{t}{\bm{K}}_{\rm B}\dot{\vec{\Psi}}_{\rm B}+\vec{\Psi}_{\rm B}^{\prime t}{\bm{G}}_{\rm B}\vec{\Psi}_{\rm B}^{\prime}+\vec{\Psi}_{\rm B}^{t}{\bm{M}}_{\rm B}\vec{\Psi}_{\rm B}+\vec{\Psi}_{\rm B}^{t}{\bm{Q}}_{\rm B}\vec{\Psi}_{\rm B}^{\prime}\,,$ (56)
where ${\bm{K}}_{\rm B}$, ${\bm{G}}_{\rm B}$, ${\bm{M}}_{\rm B}$ are $3\times 3$ symmetric matrices, ${\bm{Q}}_{\rm B}$ is an antisymmetric matrix, and
$\vec{\Psi}_{\rm B}^{t}=\left(\psi,V,\delta\phi\right)\,.$ (57)
The positivity of ${\bm{K}}_{\rm B}$ determines the no-ghost conditions. Taking the limit $l\gg 1$, they are given by
$({K_{\rm B}})_{22}=\frac{r^{4}({\cal L}_{,F}+2F{\cal L}_{,FF})^{2}}{2hN^{3/2}L{\cal L}_{,F}}>0\,,$ (59)
$({K_{\rm B}})_{22}({K_{\rm B}})_{33}-({K_{\rm B}})_{23}^{2}$
$=\frac{r^{6}({\cal L}_{,F}+2F{\cal L}_{,FF})^{2}{\cal L}_{,X}}{4h^{2}N^{2}L{\cal L}_{,F}}>0\,,$
${\rm det}\,{\bm{K}}_{\rm B}$ $=$ $\frac{M_{\rm Pl}^{2}r^{6}({\cal L}_{,F}+2F{\cal L}_{,FF})^{2}{\cal L}_{,X}}{4hN^{5/2}L^{3}{\cal L}_{,F}}>0\,.$ (60)
These inequalities are satisfied if
${\cal L}_{,X}>0,\quad{\rm and}\quad{\cal L}_{,F}>0\,,$ (61)
where the latter is the same as the no-ghost condition (51) in the odd-parity sector.
The perturbation equations of motion for $\psi$, $V$, and $\delta\phi$ follow by varying (56) with respect to these dynamical fields. We substitute the WKB-form solution
$\vec{\Psi}_{\rm B}^{t}=(\vec{\Psi}_{0}^{t})_{\rm B}e^{-i(\omega t-kr)}\,,$ (62)
into these equations, where $(\vec{\Psi}_{0}^{t})_{\rm B}=(\psi_{0},V_{0},\delta\phi_{0})$ is a constant vector. The resulting algebraic equation ${\bm{U}}_{\rm B}(\vec{\Psi}_{0})_{\rm B}=0$ has nonvanishing solutions of $(\vec{\Psi}_{0})_{\rm B}$, so long as the determinant of the $3\times 3$ matrix ${\bm{U}}_{\rm B}$ is vanishing, i.e.,
${\rm det}~{}{\bm{U}}_{\rm B}=0\,.$ (63)
The radial propagation speeds $c_{r}$ can be obtained by taking the limit $\omega r_{h}\approx kr_{h}\gg l\gg 1$ in Eq. (63) and substituting the relation $\omega=h\sqrt{N}c_{r}k$ into this equation. Then, we obtain the following three squared propagation speeds
$c_{r,\psi}^{2}=1\,,$ (64)
$c_{r,V}^{2}=1\,,$ (65)
$c_{r,\delta\phi}^{2}=1+\frac{2X[{\cal L}_{,XX}({\cal L}_{,F}+2F{\cal L}_{,FF})-2F{\cal L}_{,FX}^{2}]}{{\cal L}_{,X}({\cal L}_{,F}+2F{\cal L}_{,FF})}.$
In theories where ${\cal L}$ is a function of $F$ alone, there are two dynamical perturbations $\psi$ and $V$ that propagate with the speed of light [42]. In current theories, we have an additional scalar perturbation $\delta\phi$ whose propagation speed is different from 1. To avoid the Laplacian instability of $\delta\phi$ along the radial direction, we require that $c_{r,\delta\phi}^{2}>0$.
Taking the large multipole limit characterized by the condition $l\approx\omega r_{h}\gg kr_{h}\gg 1$ in Eq. (63), the resulting squared angular propagation speeds are given by
$c_{\Omega,\psi}^{2}=1\,,$ (67)
$c_{\Omega,V}^{2}=\frac{{\cal L}_{,F}}{{\cal L}_{,F}+2F{\cal L}_{,FF}}\,,$ (68)
$c_{\Omega,\delta\phi}^{2}=1\,.$ (69)
Here, the vector mode is decoupled from the other two modes as $K_{22}\omega^{2}+M_{22}L=0$. While both $\psi$ and $\delta\phi$ have luminal propagation speeds, $c_{\Omega,V}^{2}$ is different from 1 in theories containing nonlinear functions of $F$. The angular Laplacian instability of $V$ can be avoided if $c_{\Omega,V}^{2}>0$.
The above discussion is valid in the regime characterized by $h>0$, but we can perform a similar analysis for $h<0$. In the latter regime, the time-like and space-like properties of metric components $f~{}(=Nh)$ and $h$ are reversed compared to those in the former regime. The no-ghost conditions can be derived from the matrices ${\bm{G}}_{\rm A}$ and ${\bm{G}}_{\rm B}$ rather than ${\bm{K}}_{\rm A}$ and ${\bm{K}}_{\rm B}$. So long as ${\cal L}_{,F}>0$ and ${\cal L}_{,X}>0$, there are no ghosts in either odd-parity or even-parity sectors.
The radial and angular propagation speeds of odd- and even-parity perturbations can be derived by exploiting the WKB solutions $\vec{\Psi}_{\rm A}^{t}=(\vec{\Psi}_{0}^{t})_{\rm A}e^{-i(\omega r-kt)}$ and $\vec{\Psi}_{\rm B}^{t}=(\vec{\Psi}_{0}^{t})_{\rm B}e^{-i(\omega r-kt)}$. On using the relation $\omega=kc_{r}/(-h\sqrt{N})$ in the limit $\omega r_{h}\approx kr_{h}\gg l\gg 1$, we obtain the same values of $c_{r}^{2}$ as those given in Eqs. (54) and Eqs. (64)-(III.2.2). Taking the other limit $l\approx\omega r_{h}\gg kr_{h}\gg 1$ with the relation $\omega=c_{\Omega}l/(r\sqrt{-h})$, we find that the squared angular propagation speeds are the same as those given in Eq. (55) and Eqs. (67)-(69).
In summary, for both $h>0$ and $h<0$, the linear stability of electric SSS objects is ensured under the four conditions ${\cal L}_{,X}>0$, ${\cal L}_{,F}>0$, $c_{r,\delta\phi}^{2}>0$, and $c_{\Omega,V}^{2}>0$, where $c_{r,\delta\phi}^{2}$ and $c_{\Omega,V}^{2}$ are given, respectively, by Eqs. (III.2.2) and (68).
III.3 Stability of magnetic SSS objects
Let us proceed to the stability of magnetic SSS objects, in which case $A_{0}^{\prime}$ is vanishing. We consider the regime $h>0$, but we will briefly mention the case $h<0$ at the end of this subsection. For $q_{M}\neq 0$ and $q_{E}=0$, the second-order action (42) is decomposed into the two sectors described by the combinations of perturbations
$\vec{\Psi}_{\rm C}^{t}=\left(\chi,V\right)\,,\qquad\vec{\Psi}_{\rm D}^{t}=\left(\delta A,\psi,\delta\phi\right)\,,$ (70)
which we call sectors C and D, respectively. The vector $\vec{\Psi}_{\rm C}^{t}$ is composed of the odd-parity gravitational perturbation $\chi$ and the even-parity vector-field perturbation $V$. The vector $\vec{\Psi}_{\rm D}^{t}$ consists of the odd-parity vector-field perturbation $\delta A$, the even-parity gravitational perturbation $\psi$, and the even-parity scalar-field perturbation $\delta\phi$. Since the sectors C and D contain both odd- and even-parity modes, we deal with the total action (42) at once.
First of all, we derive the field equations of motion for $Q$ and $W$ from Eq. (42). They are used to eliminate $Q$, $W$, and their derivatives from the second-order action $\tilde{{\cal S}}^{(2)}$. Then, we vary the resulting action with respect to $H_{0}$, $H_{1}$, $\delta A_{0}$, and $\delta A_{1}$. This allows us to solve these perturbation equations for $h_{1}$, $H_{1}$, $\delta A_{0}$, and $\delta A_{1}$, so that these fields are removed from the action. After the integration by parts, the final second-order action can be expressed in the form
$\tilde{{\cal S}}^{(2)}=\int{\rm d}t{\rm d}r\left(\tilde{\cal L}_{\rm C}+\tilde{\cal L}_{\rm D}\right)\,,$ (71)
where ~$\tilde{\cal L}_{\rm C}$ and ~$\tilde{\cal L}_{\rm D}$ are the Lagrangians containing perturbations in the sectors C and D, respectively. In the following, we will address the linear stability of magnetic SSS objects for the sectors C and D, in turn.
III.3.1 Sector C
The Lagrangian ~$\tilde{{\cal L}}_{\rm C}$ is of the following form
~$\tilde{{\cal L}}_{\rm C}=\dot{\vec{\Psi}}_{\rm C}^{t}{\bm{K}}_{\rm C}\dot{\vec{\Psi}}_{\rm C}+\vec{\Psi}_{\rm C}^{\prime t}{\bm{G}}_{\rm C}\vec{\Psi}_{\rm C}^{\prime}+\vec{\Psi}_{\rm C}^{t}{\bm{M}}_{\rm C}\vec{\Psi}_{\rm C}+\vec{\Psi}_{\rm C}^{t}{\bm{Q}}_{\rm C}\vec{\Psi}_{\rm C}^{\prime}\,,$ (72)
where ${\bm{K}}_{\rm C}$, ${\bm{G}}_{\rm C}$, ${\bm{M}}_{\rm C}$ are $2\times 2$ symmetric matrices, while ${\bm{Q}}_{\rm C}$ is antisymmetric. Unlike ${\bm{K}}_{\rm A}$, the kinetic matrix ${\bm{K}}_{\rm C}$ has both diagonal and off-diagonal components. Then, the no-ghost conditions for perturbations in the sector C are given by
$(K_{\rm C})_{11}$ $=$ $\frac{M_{\rm Pl}^{2}L}{4hN^{3/2}(L-2)}>0\,,$ (73)
$(K_{\rm C})_{11}(K_{\rm C})_{22}-(K_{\rm C})_{12}^{2}$ $=$ $\frac{M_{\rm Pl}^{2}r^{4}{\cal L}_{,F}}{8h^{2}N^{3}(L-2)}>0\,.$ (74)
The first inequality (73) is automatically satisfied for $l\geq 2$, whereas the second inequality (74) holds if
${\cal L}_{,F}>0\,,$ (75)
which is the same as the no-ghost condition in the sector A.
We derive the perturbation equations for $\chi$ and $V$ from the Lagrangian (72) and substitute the WKB solution $\vec{\Psi}_{\rm C}^{t}=(\vec{\Psi}_{0}^{t})_{\rm C}\,e^{-i(\omega t-kr)}$ into them. The resulting equations are expressed in the form ${\bm{U}}_{\rm C}(\vec{\Psi}_{0})_{\rm C}=0$. Taking the limit $\omega r_{h}\approx kr_{h}\gg l\gg 1$ and using the relation $\omega=h\sqrt{N}c_{r}k$ in the determinant equation ${\rm det}\,{\bm{U}}_{\rm C}=0$, we obtain the two squared radial propagation speeds
$c_{r}^{2}=1\,,\quad{\rm for}\quad\vec{\Psi}_{\rm C}^{t}=\left(\chi,V\right)\,,$ (76)
which are both luminal. In the other limit $l\approx\omega r_{h}\gg kr_{h}\gg 1$, we substitute the relation $\omega=l\sqrt{Nh}\,c_{\Omega}/r$ into ${\rm det}\,{\bm{U}}_{\rm C}=0$. This leads to the two squared angular propagation speeds
$c_{\Omega}^{2}=1\,,\quad{\rm for}\quad\vec{\Psi}_{\rm C}^{t}=\left(\chi,V\right)\,,$ (77)
both of which are luminal as well.
III.3.2 Sector D
The Lagrangian in the sector D can be expressed in the form
~$\tilde{{\cal L}}_{\rm D}=\dot{\vec{\Psi}}_{\rm D}^{t}{\bm{K}}_{\rm D}\dot{\vec{\Psi}}_{\rm D}+\vec{\Psi}_{\rm D}^{\prime t}{\bm{G}}_{\rm D}\vec{\Psi}_{\rm D}^{\prime}+\vec{\Psi}_{\rm D}^{t}{\bm{M}}_{\rm D}\vec{\Psi}_{\rm D}+\vec{\Psi}_{\rm D}^{t}{\bm{Q}}_{\rm D}\vec{\Psi}_{\rm D}^{\prime}\,,$ (78)
where ${\bm{K}}_{\rm D}$, ${\bm{G}}_{\rm D}$, ${\bm{M}}_{\rm D}$ are $3\times 3$ symmetric matrices, whereas ${\bm{Q}}_{\rm D}$ is antisymmetric. From the kinetic matrix ${\bm{K}}_{\rm D}$, we obtain the following three no-ghost conditions
$({K_{\rm D}})_{33}$ $=$ $\frac{r^{2}{\cal L}_{,X}}{2h\sqrt{N}}>0\,,$ (79)
$({K_{\rm D}})_{22}({K_{\rm D}})_{33}$displaystyle-- $({K_{\rm D}})_{23}^{2}=\frac{M_{\rm Pl}^{2}r^{2}{\cal L}_{,X}}{2NL^{2}}>0\,,$ (80)
${\rm det}\,{\bm{K}}_{\rm D}$ $=$ $\frac{M_{\rm Pl}^{2}r^{2}{\cal L}_{,F}{\cal L}_{,X}}{4hLN^{3/2}}>0\,.$ (81)
These inequalities are satisfied if
${\cal L}_{,X}>0\,,\quad{\rm and}\quad{\cal L}_{,F}>0\,,$ (82)
which are the same as the no-ghost conditions of electric SSS objects in the even-parity sector.
Varying the Lagrangian (78) with respect to $\delta A$, $\psi$, $\delta\phi$ and using the WKB solution $\vec{\Psi}_{\rm D}^{t}=(\vec{\Psi}_{0}^{t})_{\rm D}\,e^{-i(\omega t-kr)}$, we can write the perturbation equations in the form ${\bm{U}}_{\rm D}(\vec{\Psi}_{0})_{\rm D}=0$. The squared radial propagation speeds can be obtained by taking the limit $\omega r_{h}\approx kr_{h}\gg l\gg 1$ in the determinant equation ${\rm det}\,{\bm{U}}_{\rm D}=0$, leading to
$c_{r,\delta A}^{2}=1\,,$ (83)
$c_{r,\psi}^{2}=1\,,$ (84)
$c_{r,\delta\phi}^{2}=1+\frac{2X{\cal L}_{,XX}}{{\cal L}_{,X}}\,.$ (85)
In theories with ${\cal L}={\cal L}(F)$, the dynamical perturbations $\delta A$ and $\psi$ propagate with the speed of light, as consistent with the result in Ref. [42]. In current theories, the additional scalar DOF $\delta\phi$ has the propagation speed different from 1. As we can compare with Eq. (III.2.2), the electric SSS object has a different value of $c_{r,\delta\phi}^{2}$ in comparison to Eq. (85). In theories with ${\cal L}_{,FX}=0$, they are identical to each other.
Taking the other limit $l\approx\omega r_{h}\gg kr_{h}\gg 1$ in the determinant equation ${\rm det}\,{\bm{U}}_{\rm D}=0$, we obtain the following squared angular propagation speeds
$c_{\Omega,\delta A}^{2}=1+\frac{2F{\cal L}_{,FF}}{{\cal L}_{,F}}\,,$ (86)
$c_{\Omega,\psi}^{2}=1\,,$ (87)
$c_{\Omega,\delta\phi}^{2}=1\,.$ (88)
If ${\cal L}$ contains nonlinear functions of $F$, the propagation speed of $\delta A$ is different from 1. The expression of $c_{\Omega,\delta A}^{2}$ coincides with the one derived in Ref. [42] for theories with ${\cal L}={\cal L}(F)$.
We have also studied the case $h<0$ and obtained the same no-ghost conditions and radial/angular propagation speeds as those derived for $h>0$. In summary, the linear stability of magnetic SSS objects requires that the four conditions ${\cal L}_{,X}>0$, ${\cal L}_{,F}>0$, $c_{r,\delta\phi}^{2}>0$, and $c_{\Omega,\delta A}^{2}>0$ are satisfied, where $c_{r,\delta\phi}^{2}$ and $c_{\Omega,\delta A}^{2}$ are given, respectively by Eqs. (85) and (86).
IV k-essence theories with NED
As a first example of the possible realization of nonsingular SSS objects, we will discuss the case of NED in the presence of a k-essence scalar field. The k-essence Lagrangian of the form $K(\phi,X)$ was originally introduced in the context of inflation and dark energy [53, 54, 55]. Now, we consider the following Lagrangian
=~${\cal L}=\tilde{{\cal L}}(F)+K(\phi,X)\,,$ (89)
where $\tilde{{\cal L}}$ is a function of $F$ alone, and $K$ depends on $\phi$ and $X$. Since the electromagnetic field is not directly coupled to the scalar field, we have that ${\cal L}_{,F}=\tilde{{\cal L}}_{,F}$ and ${\cal L}_{,FX}=0$. We note that nonsingular magnetic black-bounce solutions were recently studied in Einstein gravity with the Lagrangian (89) [56]. In the following, we will study the stability of electric and magnetic SSS objects in turn.
IV.1 Electric case
For $q_{E}\neq 0$ and $q_{M}=0$, the squared propagation speeds (III.2.2) and (68) in the even-parity sector yield
$c_{r,\delta\phi}^{2}$ $=$ $1+\frac{2XK_{,XX}}{K_{,X}}\,,$ (90)
$c_{\Omega,V}^{2}$ $=$ ~$\frac{\tilde{{\cal L}}_{,F}}{\tilde{{\cal L}}_{,F}+2F\tilde{{\cal L}}_{,FF}}\,,$ (91)
whereas all the other dynamical perturbations propagate with the speed of light. We can compute ~$\tilde{{\cal L}}_{,FF}$ by taking the $r$ derivatives of ~$\tilde{{\cal L}}_{,F}=q_{E}\sqrt{N}/(r^{2}A_{0}^{\prime})$ and $F=A_{0}^{\prime 2}/(2N)$, as ~$\tilde{{\cal L}}_{,FF}=\tilde{{\cal L}}_{,F}^{\prime}(r)/F^{\prime}(r)$. Then, Eq. (91) reduces to
$c_{\Omega,V}^{2}=\frac{r(A_{0}^{\prime}N^{\prime}-2A_{0}^{\prime\prime}N)}{4A_{0}^{\prime}N}\,.$ (92)
Since $A_{0}^{\prime}$ is given by Eq. (21), we can express $c_{\Omega,V}^{2}$ in terms of $h$, $N$, and their $r$ derivatives, as
$c_{\Omega,V}^{2}$ $=$ $-r(2N^{3}r^{2}h^{\prime\prime\prime}+2hr^{2}N^{2}N^{\prime\prime\prime}+3r^{2}N^{2}N^{\prime}h^{\prime\prime}+5r^{2}h^{\prime}N^{2}N^{\prime\prime}-4hr^{2}NN^{\prime}N^{\prime\prime}-4h^{\prime}r^{2}NN^{\prime 2}+2hr^{2}N^{\prime 3}$ (93)
$+4N^{3}rh^{\prime\prime}+6hrN^{2}N^{\prime\prime}+8h^{\prime}rN^{2}N^{\prime}-4hrNN^{\prime 2}-4h^{\prime}N^{3}+2hN^{2}N^{\prime})$
$/[{2N(2N^{2}r^{2}h^{\prime\prime}+2hr^{2}NN^{\prime\prime}+3h^{\prime}r^{2}NN^{\prime}-hr^{2}N^{\prime 2}+2hrNN^{\prime}-4hN^{2}+4N^{2})}]\,.$
Using the expansions (17) and (18) of $h$ and $N$ around $r=0$, we obtain
$c_{\Omega,V}^{2}=-1-\frac{4N_{0}h_{3}+9N_{3}}{8N_{2}}r+{\cal O}(r^{2})\,.$ (94)
Nonsingular BHs studied in the literature typically have the properties $h_{3}=0$ and $N_{3}=0$ [2, 25, 4, 5]. In such cases, the expansion of $c_{\Omega,V}^{2}$ around $r=0$ leads to
$c_{\Omega,V}^{2}$ $=$ $-1+\frac{3N_{2}^{2}-N_{0}(5h_{2}N_{2}+8N_{4})-5h_{4}N_{0}^{2}}{2N_{0}N_{2}}r^{2}$ (95)
$+{\cal O}(r^{3})\,.$
Since the leading-order contributions to $c_{\Omega,V}^{2}$ are negative for both the cases (94) and (95), the vector-field perturbation $V$ is subject to Laplacian instability in the angular direction. We note that NED without the scalar field corresponds to $N(r)=1$ for all $r$. In this case, Eq. (93) reduces to the value of $c_{\Omega,V}^{2}$ derived in Ref. [42]. For $N(r)=1$, the leading-order term of $c_{\Omega,V}^{2}$ is also negative.
Since the vector-field perturbation is coupled to the gravitational perturbation, the Laplacian instability of $V$ leads to the enhancement of $\psi$ along the angular direction. As studied in Ref. [42], the typical time scale of instability can be estimated as $t_{\rm ins}\simeq r/(\sqrt{-c_{\Omega}^{2}}\,l)$. For $l\gg 1$, $t_{\rm ins}$ is infinitely small. Due to this rapid growth of even-parity perturbations around $r=0$, the line element of nonsingular electric SSS objects cannot be sustained in a steady state. We note that the choice of the scalar-field Lagrangian $K(\phi,X)$ does not affect the discussion given above. In other words, no matter how we choose the functional forms of $K(\phi,X)$, the angular instabilities of $V$ and $\psi$ are inevitable for electric SSS objects.
IV.2 Magnetic case
For the magnetic case, the squared propagation speeds (85) and (86) in the sector D reduce, respectively, to
$c_{r,\delta\phi}^{2}$ $=$ $1+\frac{2XK_{,XX}}{K_{,X}}\,,$ (96)
$c_{\Omega,\delta A}^{2}$ $=$ ~$\frac{\tilde{{\cal L}}_{,F}+2F\tilde{{\cal L}}_{,FF}}{\tilde{{\cal L}}_{,F}}\,,$ (97)
while all the other dynamical perturbations have luminal propagation speeds. From Eq. (26), we have
~$\tilde{\cal L}(F)+K(\phi,X)=\frac{M_{\rm Pl}^{2}}{r^{2}}\left(rh^{\prime}+h-1\right)\,,$ (98)
with $F=-q_{M}^{2}/(2r^{4})$. Taking the $r$ derivative of Eq. (98) and exploiting Eqs. (9) and (11), we find that ~$\tilde{\cal L}_{,F}$ is written as
~$\tilde{\cal L}_{,F}$ $=$ $M_{\rm Pl}^{2}r^{2}[2N^{2}(r^{2}h^{\prime\prime}-2h+2)+2rhN(rN^{\prime\prime}+N^{\prime})$ (99)
$+3r^{2}h^{\prime}N^{\prime}N-r^{2}hN^{\prime 2}]/(4q_{M}^{2}N^{2})\,.$
Differentiating this equation with respect to $r$, we can express ~$\tilde{\cal L}_{,FF}$ in terms of $N$, $h$, and their $r$ derivatives. Then, it follows that $c_{\Omega,\delta A}^{2}$ is completely identical to $c_{\Omega,V}^{2}$ for the electric configuration given by Eq. (93).
Using the expansions of $h$ and $N$ around $r=0$, the leading-order contribution to $c_{\Omega,\delta A}^{2}$ is $-1$ and hence the odd-parity vector-field perturbation $\delta A$ is subject to angular Laplacian instability. This leads to the enhancement of the gravitational perturbation $\psi$. Then, the nonsingular magnetic SSS object cannot be present as a stable configuration.
In summary, for theories with the Lagrangian (89), we have shown that all the nonsingular electric and magnetic SSS objects are excluded by angular Laplacian instabilities arising from vector-field perturbations. This includes nonsingular BHs constructed from the Lagrangian (89), which extends our previous results found for NED [42].
V Theories with ${\cal L}=X+\mu(\phi)F^{n}$
In theories where the Lagrangian ${\cal L}$ contains the nonlinear dependence of $F$, ${\cal L}_{,FF}$ does not vanish in Eq. (68) or Eq. (86). As we showed in Sec. IV, this results in the negative values of $c_{\Omega,V}^{2}$ or $c_{\Omega,\delta A}^{2}$ around $r=0$. If we consider theories in which ${\cal L}$ contains only a linear term in $F$, it is possible to avoid the problem of angular Laplacian instabilities. We also note that the nonlinear dependence of $X$ in ${\cal L}$ leads to the deviation of $c_{r,\delta\phi}^{2}$ from 1. The linear term in $X$ without a direct coupling to $F$ results in the value $c_{r,\delta\phi}^{2}=1$ for both electric and magnetic cases.
In this section, we consider theories given by the Lagrangian
${\cal L}=X+\mu(\phi)F^{n}\,,$ (100)
where $\mu$ is a function of $\phi$, and $n$ is an integer. Einstein-Maxwell-scalar theories correspond to the particular power $n=1$. A dilaton field in string theory has an exponential coupling $\mu(\phi)=\mu_{0}e^{-\lambda\phi}$ [57]. In such Einstein-Maxwell-dilaton theories, it is known that there is an exact hairy BH solution with a singularity at $r=0$ [58, 59]. If $\mu$ contains even power-law functions of $\phi$, tachyonic instability of the RN branch can give rise to scalarized charged BH solutions with curvature singularities at $r=0$ [60, 61, 62, 63]. In this work, we would like to explore whether stable nonsingular BHs and compact objects can be present in Einstein-Maxwell-scalar theories for general power $n$.
The absence of nonlinear terms in $X$ leads to the luminal propagation of the scalar field, i.e.,
$c_{r,\delta\phi}^{2}=1\,,$ (101)
for both electric and magnetic configurations. From Eqs. (68) and (86), we have
$c_{\Omega,V}^{2}$ $=$ $\frac{1}{2n-1},\qquad{\rm for~{}the~{}electric~{}case}\,,$ (102)
$c_{\Omega,\delta A}^{2}$ $=$ $2n-1,\qquad\,{\rm for~{}the~{}magnetic~{}case}\,.$ (103)
Then, the angular Laplacian instability is absent if
$n>\frac{1}{2}\,.$ (104)
One of the no-ghost conditions ${\cal L}_{,X}>0$ is trivially satisfied. On the other hand, the other no-ghost condition ${\cal L}_{,F}>0$ is given by
${\cal L}_{,F}=n\mu(\phi)\left(\frac{A_{0}^{\prime 2}}{2N}-\frac{q_{M}^{2}}{2r^{4}}\right)^{n-1}>0\,.$ (105)
For integer-odd values of $n$ in the range (104), this inequality translates to $\mu(\phi)>0$. In the following, we will study the two cases: (i) $q_{E}\neq 0$, $q_{M}=0$, and (ii) $q_{M}\neq 0$, $q_{E}=0$, in turn. In case (ii), if $n$ is not an integer, we should deal with $F^{n}$ in the Lagrangian, as $(F^{2})^{m}$ with $n=2m$, due to the negativity of $F$.
V.1 Electric case
For $q_{E}\neq 0$ and $q_{M}=0$, Eqs. (12) and (14) give
$\mu(\phi)$ $=$ $\frac{q_{E}\sqrt{N}}{nr^{2}A_{0}^{\prime}F^{n-1}}\,,$ (106)
$\phi^{\prime}$ $=$ $M_{\rm Pl}\sqrt{\frac{N^{\prime}}{rN}}\,,$ (107)
where we have chosen the branch $\phi^{\prime}>0$ without loss of generality. For the realization of the solution (107), we require that $N^{\prime}>0$. From Eqs. (13) and (100), we have
$-\frac{1}{2}h\phi^{\prime 2}+\mu(\phi)\left(\frac{A_{0}^{\prime 2}}{2N}\right)^{n}=\frac{M_{\rm Pl}^{2}}{r^{2}}\left(rh^{\prime}+h-1\right)+\frac{q_{E}A_{0}^{\prime}}{\sqrt{N}r^{2}}\,.$ (108)
Substituting Eqs. (21), (106), (107) into Eq. (108), we find that $h(r)$ and $N(r)$ are related to each other, as
$2[(2n-1)r^{2}h^{\prime\prime}+4nrh^{\prime}+2h-2]N^{2}-(2n-1)r^{2}hN^{\prime 2}$
$+[3(2n-1)r^{2}h^{\prime}+2(4n-1)rh]NN^{\prime}$
$+2(2n-1)r^{2}hNN^{\prime\prime}=0\,.$ (109)
By using this relation, we can express $A_{0}^{\prime}$ in Eq. (21) in a simpler form
$\frac{q_{E}A_{0}^{\prime}}{M_{\rm Pl}^{2}}=-\frac{n\left[2\left(rh^{\prime}+h-1\right)N+rhN^{\prime}\right]}{\sqrt{N}\,\left(2n-1\right)}\,.$ (110)
Since $n>1/2$, the integrated solution to Eq. (109) is expressed in the form
$h(r)$ $=$ $\frac{2\int_{0}^{r}r_{1}^{-\frac{4(n-1)}{2n-1}}\sqrt{N(r_{1})}\bigl{(}\int_{0}^{r_{1}}\sqrt{N(r_{2})}{\rm d}r_{2}\bigr{)}{\rm d}r_{1}}{(2n-1)r^{\frac{2}{2n-1}}N(r)}$ (111)
$+\frac{c_{1}}{r^{\frac{2}{2n-1}}N(r)}+\frac{c_{2}\int_{0}^{r}r_{1}^{-\frac{4(n-1)}{2n-1}}\sqrt{N(r_{1})}{\rm d}r_{1}}{(2n-1)r^{\frac{2}{2n-1}}N(r)}\,,$
where $c_{1}$ and $c_{2}$ are integration constants. We recall that $N(r)$ is expanded as Eq. (18) around $r=0$. To avoid the divergence of $h(r)$ at $r=0$, we require that $c_{1}=0=c_{2}$. Then, the solution (111) reduces to
$h(r)=\frac{2\int_{0}^{r}r_{1}^{-\frac{4(n-1)}{2n-1}}\sqrt{N(r_{1})}\bigl{(}\int_{0}^{r_{1}}\sqrt{N(r_{2})}{\rm d}r_{2}\bigr{)}{\rm d}r_{1}}{(2n-1)r^{\frac{2}{2n-1}}N(r)}.$ (112)
Under the linear stability condition (104), $h(r)$ is always positive at any distance $r\geq 0$. Therefore, we do not have the nonsingular BH configuration with an apparent horizon.
On using the expansion (18) of $N(r)$ around $r=0$, we find that Eq. (112) can be expanded as
$h(r)=1-\frac{3n-1}{3n}\frac{N_{2}}{N_{0}}r^{2}+{\cal O}(r^{3})\,,$ (113)
whose dependence is analogous to the standard boundary condition of metrics of stars. However, we need to make sure that the radial derivatives of $A_{0}$ and $\phi$ are vanishing at $r=0$ for the realization of regular SSS objects. From Eq. (22), the leading-order term of $A_{0}^{\prime}(r)$ is proportional to $r^{2}$ and hence $A_{0}^{\prime}(r)\to 0$ as $r\to 0$. Applying the expansion (18) to Eq. (107) for $N_{2}\neq 0$, it follows that
$\frac{\phi^{\prime}(r)}{M_{\rm Pl}}=\sqrt{\frac{2N_{2}}{N_{0}}}+\frac{3}{4N_{2}}\sqrt{\frac{2N_{2}}{N_{0}}}N_{3}\,r+{\cal O}(r^{2})\,.$ (114)
As $r\to 0$, $\phi^{\prime}(r)$ approaches a nonvanishing constant $M_{\rm Pl}\sqrt{2N_{2}/N_{0}}$. This can lead to a cusp-like structure around the origin, whose property should be incompatible with the SSS background. To avoid this behavior, we need to impose the condition $N_{2}=0$. Moreover, the realization of the regular solution $\phi^{\prime}\propto r$ around $r=0$ requires that $N_{3}=0$. Then, so long as the expansion of $N(r)$ around $r=0$ is given by
$N(r)=N_{0}+N_{4}r^{4}+{\cal O}(r^{5})\,,$ (115)
the scalar-field derivative and the kinetic term have the following behavior
$\phi^{\prime}(r)$ $=2M_{\rm Pl}\sqrt{\frac{N_{4}}{N_{0}}}r+{\cal O}(r^{2})\,,$ (116)
$X(r)$ $=-2M_{\rm Pl}^{2}\frac{N_{4}}{N_{0}}r^{2}+{\cal O}(r^{3})\,.$ (117)
From Eq. (116), we require that $N_{4}>0$. We also have the following expansions
$h(r)$ $=$ $1-\frac{20n-8}{20n-5}\,\frac{N_{4}}{N_{0}}\,r^{4}+\mathcal{O}(r^{5})\,,$ (118)
$A_{0}^{\prime}(r)$ $=$ $\frac{12M_{\rm Pl}^{2}nN_{4}}{q_{E}\sqrt{N_{0}}(4n-1)}r^{4}+{\cal O}(r^{5})\,,$ (119)
so that $F=\mathcal{O}(r^{8})$. We notice that $h(r)<1$ under the inequalities $n>1/2$, $N_{0}>0$, and $N_{4}>0$. Thus, $N(r)$ is constrained to be of the form (115) to realize the regular behavior $\phi^{\prime}(r)\propto r$ around the origin.
One of the examples for $N(r)$ that has the property (115) is given by
$N(r)=\left(\frac{r^{4}+\sqrt{N_{0}}r_{0}^{4}}{r^{4}+r_{0}^{4}}\right)^{2}\,,$ (120)
where $N_{0}$ and $r_{0}$ are positive constants. Around $r=0$, this function is expanded as
$N(r)=N_{0}+\frac{2\sqrt{N_{0}}(1-\sqrt{N_{0}})}{r_{0}^{4}}r^{4}+{\cal O}(r^{8})\,,$ (121)
and hence $N_{4}=2\sqrt{N_{0}}(1-\sqrt{N_{0}})/r_{0}^{4}$. Since the inequality $N_{4}>0$ holds, $N_{0}$ should be in the range
$0<N_{0}<1\,.$ (122)
At spatial infinity, the function (120) approaches 1 with a correction of order $r^{-4}$.
Let us pause for a moment to reflect on the reasoning behind the choice (120). In the standard approach, one typically specifies the function $\mu(\phi)$ and then searches for solutions by imposing boundary conditions that are consistent with the given choice of $\mu(\phi)$. In contrast, we adopt a different strategy here. We prescribe the form of $N(r)$ and determine the corresponding $\phi(r)$ and $\mu(r)$ by integrating Eqs. (107) and (110), together with the relation (108).
As $\phi$ evolves from $\phi_{0}$ (its value at the origin) to $\phi_{\infty}$ (its value at infinity), we will demonstrate that it is, in principle, possible to reconstruct $\mu(\phi)$ for $\phi_{0}<\phi<\phi_{\infty}$, assuming the branch where $\phi^{\prime}>0$. Once $\mu(\phi)$ is obtained–at least over this range of $\phi$–we can revert to the conventional approach of solving for configurations compatible with the reconstructed $\mu(\phi)$.
In general, the set of solutions corresponding to this $\mu(\phi)$, if it is neither empty nor a singleton, will be distinguished by their respective values of mass and charge of the object. [^2] This would mean that, for a fixed $\mu(\phi)$, there would be more families of profiles, e.g., for $N(r)$ and $\phi(r)$. The variation in these parameters may result in different forms of $N(r)$, [^3] Notice, however, that for the RN solution in GR, although the mass and charge vary, the function $N$ is unity over the whole manifold. For perfect fluids, we instead have $N\neq 1$, when a non-negligible pressure is present. which must satisfy the boundary conditions $N^{\prime}(r\to 0)\propto r^{3}$ at the origin and $N\to 1$ at infinity. [^4] In addition, we require the scalar-field dependence $\phi^{\prime}(r\to 0)\propto r$ as well as all the other boundary conditions that define regular objects with asymptotically flat spacetime. Exploring this second avenue, checking the properties of solutions for a fixed $\mu(\phi)$ is an interesting prospect. Still, we believe that its investigation lies beyond the scope of this work and should be addressed in a future study.
For $n$ in the range $n>1/2$ with the choice (120), Eq. (112) shows that $h(r)$ also approaches 1 as $r\to\infty$. Then, the background metric components satisfy the condition for asymptotic flatness. At large distances ($r\gg r_{0}$), the differential Eq. (109) is approximately given by
$(2n-1)h^{\prime\prime}\simeq-\frac{4n}{r}h^{\prime}-\frac{2}{r^{2}}h+\frac{2}{r^{2}}\,,$ (123)
where we have kept the most dominant $r$-dependent contributions in the coefficients of $h^{\prime}$ and $h$. For $n>1/2$, we can integrate Eq. (123) to give
$h(r)=1+\frac{c_{1}}{r}+c_{2}r^{-\frac{2}{2n-1}}\,,\qquad{\rm for}\quad n\neq\frac{3}{2}\,,$ (124)
and
$h=1+\frac{c_{1}}{r}+c_{2}\frac{\ln\,r}{r}\,,\qquad{\rm for}\quad n=\frac{3}{2}\,,$ (125)
where $c_{1}$ and $c_{2}$ are constants. Up to the next-to-leading order terms to $h(r)$, we can classify the large-distance behavior of $h(r)$, as
$h(r)$ $\simeq$ $1+\frac{c_{1}}{r}\,,\qquad{\rm for}\quad\frac{1}{2}<n<\frac{3}{2}\,,$ (126)
$h(r)$ $\simeq$ $1+c_{2}\frac{\ln r}{r}\,,\qquad{\rm for}\quad n=\frac{3}{2}\,,$ (127)
$h(r)$ $\simeq$ $1+c_{2}r^{-\frac{2}{2n-1}}\,,\qquad{\rm for}\quad n>\frac{3}{2}\,.$ (128)
In the last case (128), the term $r^{-\frac{2}{2n-1}}$ decreases slowly relative to $r^{-1}$.
The Arnowitt-Deser-Misner (ADM) mass of SSS objects is defined by
$M(r)\equiv\lim_{r\to\infty}\frac{r}{2G}\,(1-h)\,.$ (129)
where $G$ is the gravitational constant. From Eqs. (127) and (128), for $n\geq 3/2$, the quantity $r(1-h)$ increases at large distances. In this case, we do not have compact SSS objects. For $n$ in the range
$\frac{1}{2}<n<\frac{3}{2}\,,$ (130)
the metric component behaves as Eq. (126) and hence $M(r)$ approaches a constant value $-c_{1}/(2G)$. This is the case in which nonsingular SSS objects satisfy the condition for compactness.
Applying Eq. (120) to Eq. (107), the scalar-field derivative has the following behavior in the regime $r\gg r_{0}$:
$\phi^{\prime}(r)=\frac{2M_{\rm Pl}r_{0}^{2}\sqrt{2(1-\sqrt{N_{0}})}}{r^{3}}+{\cal O}(r^{-7})\,.$ (131)
We note that the dependence $\phi^{\prime}(r)\propto r^{-3}$ is an outcome of the particular choice (120). If we consider $N(r)$ with the large-distance behavior $N(r)=1+{\cal O}(r^{-2})$, then we have the dependence $\phi^{\prime}(r)\propto r^{-2}$. Substituting the solution (124) of $h(r)$ into Eq. (110), we find
$\frac{q_{E}A_{0}^{\prime}}{M_{\rm Pl}^{2}}=-\frac{2c_{2}\left(2n-3\right)n}{\left(2n-1\right)^{2}r^{\frac{2}{2n-1}}}+\mathcal{O}(r^{-4})\,,$ (132)
where the term proportional to $r^{-{\frac{2}{2n-1}}}$ dominates over $r^{-4}$ for $23/4<n<3/$ 2. As we will see later in Fig. 5, the allowed values of $n$ that are consistent with no-ghost conditions are indeed larger than $3/4$.
Figure 1: Metric component $h(r)$ versus $r/r_{0}$ for electric SSS objects present in theories given by the Lagrangian (100). We choose $N(r)$ of the form (120) with $N_{0}=0.2$. Each case corresponds to (a) $n=1/2$, (b) $n=1$, (c) $n=2$, (d) $n=5$, and (e) $n\gg 1$. For $n$ in the range $n>1/2$, the theoretical lines of $h(r)$ are between (a) and (e), so that $h(r)$ is always positive at any distance $r$.
Figure 2: ADM mass $M$ (unit of $G=1$) versus $r/r_{0}$ for electric SSS objects present in theories given by the Lagrangian (100). We choose $N(r)$ to be (120) with $N_{0}=0.2$. From bottom to top, each line corresponds to $n=1/2,4/5,1,5/4,3/2,2,4$. For $n<3/2$, $M$ asymptotically approaches a constant, while, for $n\geq 3/2$, $M$ grows in the regime $r\gg r_{0}$.
In the limit $n\to 1/2$, both $h^{\prime\prime}(r)$ and $N^{\prime\prime}(r)$ vanish in Eq. (109). In this case, the integrated solution to Eq. (109) is given by
$h(r)=\frac{1}{r\sqrt{N(r)}}\left(\int_{0}^{r}\sqrt{N(r_{1})}\,{\rm d}r_{1}+c_{1}\right)\,,$ (133)
where the constant $c_{1}$ should be 0 to avoid the divergence of $h(r)$ at $r=0$. Then, Eq. (133) reduces to
$h_{n\to 1/2}(r)=\frac{1}{r\sqrt{N(r)}}\int_{0}^{r}\sqrt{N(r_{1})}\,{\rm d}r_{1}\,,$ (134)
which is positive at any distance $r\geq 0$.
In the limit that $n\gg 1$, Eq. (109) yields
$h^{\prime\prime}=\frac{r(hN^{\prime 2}-3NN^{\prime}h^{\prime}-2hNN^{\prime\prime})-4N(Nh^{\prime}+hN^{\prime})}{2rN^{2}}.$ (135)
The solution to this equation can be written as
$h(r)=\frac{c_{1}}{N(r)}+\frac{c_{2}}{N(r)}\int_{0}^{r}\frac{\sqrt{N(r_{1})}}{r_{1}^{2}}{\rm d}r_{1}\,.$ (136)
To match this with the expansions (17) and (18) around $r=0$, the integration constants are fixed to be $c_{1}=N_{0}$ and $c_{2}=0$. Then, we obtain
$h_{n\gg 1}(r)=\frac{N_{0}}{N(r)}\,,$ (137)
which is always positive. At spatial infinity, $h_{n\gg 1}(r)$ approaches $N_{0}$, so the condition for asymptotic flatness is not satisfied unless $N_{0}=1$. For finite values of $n$ in the range $n>1/2$, however, we have already seen that $h(r)$ approaches 1 as $r\to\infty$.
In Fig. 1, we plot $h$ versus $r/r_{0}$ for $n=1/2,1,2,5$ and $n\gg 1$, with the choice $N_{0}=0.2$ in Eq. (120). In each case, we integrate Eq. (109) outward from $r=0$ by using the boundary condition (118). Except for the limit $n\gg 1$, the initial decrease of $h(r)$, which is characterized by the solution (118), changes to its growth toward the asymptotic value $h(r)\to 1$. In the limit $n\gg 1$, i.e., curve (e) in Fig. 1, $h(r)$ monotonically decreases toward the asymptotic value $N_{0}$, with $h(r)>0$ for any distance $r$. Under the linear stability condition $n>1/2$, the theoretical curves of $h(r)$ lie between (a) and (e) in Fig. 1. In this region, as we showed analytically, we have $h(r)>0$ at any distance $r$. Thus, so long as the linear stability conditions are satisfied, we do not realize nonsingular BHs where $h(r)$ becomes negative for some range of $r$.
In Fig. 2, we plot the ADM mass (129) for seven different values of $n$, with $N_{0}=0.2$. The analytic estimation (126) of $h(r)$ shows that $M$ asymptotically approaches constants for 1/$21/2<n<3/$ 2, whose property can be confirmed in Fig. 2. For $n=3/2$, the ADM mass exhibits the logarithmic growth $M\propto\ln r$ due to the property of Eq. (127). For $n>3/2$, Eq. (128) gives the analytic dependence $M\propto r^{\frac{2n-3}{2n-1}}$. In Fig. 2, we can confirm this power-law growth of $M$ for $n>3/2$ at large distances. Thus, so long as 1/$21/2<n<3/$ 2, the electric SSS configuration without the apparent horizon can be interpreted as a compact object.
Figure 3: We plot $r_{0}\phi^{\prime}/M_{\rm Pl}$, $q_{E}A_{0}^{\prime}/M_{\rm Pl}^{2}$, $f$, and $h$ versus $r/r_{0}$ for the electric BH in theories given by the Lagrangian (100) with $n=1$. We choose $N(r)$ to be (120) with $N_{0}=0.2$. We observe that $\phi^{\prime}$ and $A_{0}^{\prime}$ approach 0 as $r\to 0$. The two metric components $f$ and $h$ are different around $r=0$, but both converge to 1 at spatial infinity.
Figure 4: The coupling $\mu$ versus $r/r_{0}$ for the electric SSS object with the choice $N_{0}=0.2$ in Eq. (120). We choose five different powers: $n=0.6,0.7,0.9,1,2$. For $n>0.887$, $\mu$ is positive at any distance $r$.
Figure 5: The allowed region of parameter space for electric compact objects, represented by the light-blue shaded area. The lower boundary corresponds to the values of $n_{\rm min}$, below which the no-ghost condition is violated for a given $N_{0}$ in the range $10<N_{0}<$ 1. The upper boundary, represented by $n=3/2$, marks the limit beyond which the solutions no longer describe compact objects. Thus, the allowed range for $n$ is $n_{\rm min}\leq n<3/2$. The curve of $n_{\rm min}(N_{0})$ is obtained by interpolating over a set of 99 numerically computed values of $n_{\rm min}$ for $N_{0}$ in the range $0.01\leq N_{0}\leq 0.99$.
In Fig. 3, we show $r_{0}\phi^{\prime}/M_{\rm Pl}$, $q_{E}A_{0}^{\prime}/M_{\rm Pl}^{2}$, $f$, and $h$ as functions of $r/r_{0}$ for $n=1$ with $N_{0}=0.2$ in Eq. (120). As estimated from Eqs. (116) and (119), numerical results show the properties $\phi^{\prime}\propto r$ and $A_{0}^{\prime}\propto r^{4}$ around $r=0$. According to the analytic estimations given in Eqs. (131) and (132), $\phi^{\prime}$ and $A_{0}^{\prime}$ should have the large-distance behaviors $\phi^{\prime}\propto r^{-3}$ and $A_{0}^{\prime}\propto r^{-2}$ for $n=1$. Indeed, after reaching the maximum values at intermediate distances, we confirm that $\phi^{\prime}$ and $A_{0}^{\prime}$ approach these asymptotic solutions. As we showed analytically, the asymptotic behavior of $A_{0}^{\prime}$ is different depending on the power $n$.
The metric component $f=Nh$ differs from $h$ by the factor $N$, whose difference is largest at $r=0$ by the factor $N_{0}=0.2$. Since $N$ grows from $N_{0}$ to 1 for increasing $r$, $f$ approaches $h$ at large distances and they finally converge to the asymptotic values $f\to 1$ and $h\to 1$. The analytic estimation (124) shows that $h(r)$ has the large-distance behavior $h(r)=1+c_{1}/r+c_{2}/r^{2}$ for $n=1$. From Fig. 3, we find that $h(r)<1$ and hence $c_{1}$ is negative. As we see in Fig. 2 for $n=1$, the ADM mass $M$ approaches a constant positive value $-c_{1}/(2G)$ at spatial infinity.
Let us study the bounds on $n$ derived by the no-ghost condition (105), i.e., ${\cal L}_{,F}=n\mu(\phi)F^{n-1}>0$, where $F=A_{0}^{\prime 2}/(2N)$. Using Eqs. (21), (108) and (109), this condition translates to
${\cal L}_{,F}=\frac{q_{E}^{2}(1-2n)N}{M_{\rm Pl}^{2}nr^{2}[rhN^{\prime}+2(rh^{\prime}+h-1)N]}>0\,.$ (138)
Since we are now considering the case 1/$21/2<n<3/$ 2 with $F>0$, the no-ghost condition for electric compact objects is equivalent to $\mu>0$, i.e.,
$\mu$ $=$ $\frac{2^{n-1}M_{\rm Pl}^{2}[2N(1-h-rh^{\prime})-rhN^{\prime}]}{(2n-1)r^{2}N}$ (139)
×$\times\left[\frac{q_{E}^{2}(2n-1)^{2}N^{2}}{M_{\rm Pl}^{4}n^{2}\{2(rh^{\prime}+h-1)N+rhN^{\prime}\}^{2}}\right]^{n}>0\,.$
For given values of $N_{0}$, we study how the no-ghost condition $\mu>0$ depends on the parameter $n$ within the allowed range 1/$21/2<n<3/$ 2. Our numerical analysis reveals the existence of a minimum value of $n$ for each $N_{0}$, denoted as $n_{\rm min}$, above which $\mu>0$ for all $r$. Conversely, if $n<n_{\rm min}$, $\mu$ becomes negative at some finite distance $r$, so these values of $n$ are excluded.
In Fig. 4, we plot $\mu$ versus $r/r_{0}$ with the choice $N_{0}=0.2$ for five different values of $n$. In this case, if $n<0.887$, the coupling enters the region $\mu<0$ for some values of $r$. For $n>n_{\rm min}=0.887$, we observe in Fig. 4 that $\mu>0$ at any distance $r$. For $N_{0}$ in the range $10<N_{0}<$ 1, we numerically obtain the minimum values of $n$ and plot $n_{\rm min}$ versus $N_{0}$ in Fig. 5. We find that $n_{\rm min}$ mildly increases as a function of $N_{0}$ and approaches 1 for $N_{0}\to 1$. Note that $n_{\rm min}$ is always larger than $1/2$ for any values of $N_{0}$. So long as $n$ is in the range
$n_{\rm min}\leq n<\frac{3}{2}\,,$ (140)
there exist horizonless compact objects with neither ghosts nor Laplacian instabilities.
The integrated solution to Eq. (107) is given by
$\phi(r)=\phi_{\infty}-\int_{r}^{\infty}\sqrt{\frac{N^{\prime}}{\tilde{r}N}}{\rm d}\tilde{r}\,,$ (141)
where $\phi_{\infty}$ is the value of $\phi$ at $r\to\infty$, a value that could be matched to the cosmological value of the field. Since we are considering the branch $\phi^{\prime}(r)>0$, the scalar field increases as a function of $r$. We can invert the relation $\phi=\phi(r)$ to write $r=\tilde{r}(\phi)$. Since the coupling (139) depends on $r$, i.e., $\mu=\tilde{\mu}(r)$, we can express it in the form $\mu(\phi)=\tilde{\mu}(\tilde{r}(\phi))$. In other words, the functional form of $\mu(\phi)$ is determined to realize a desired expression of $N(r)$ consistent with the boundary conditions at $r=0$ and at spatial infinity.
For example, let us consider the case $n=1$, with $N(r)$ given by Eq. (120). Around $r=0$, the coupling (139) has the dependence $\mu\propto r^{-6}$. In this regime, the scalar field behaves as $\phi=\phi_{0}+M_{\rm Pl}\sqrt{N_{4}/N_{0}}\,r^{2}+{\cal O}(r^{3})$ from Eq. (116), where $\phi_{0}=\phi(r=0)$, so that $\mu(\phi)\propto(\phi-\phi_{0})^{-3}$. Even though $\mu$ is divergent as $r\to 0$, the product $\mu(\phi)F$ in the Lagrangian approaches 0 due to the dependence $F\propto r^{8}$. At large distances, integrating Eq. (131) leads to the solution $\phi=\phi_{\infty}-M_{\rm Pl}^{2}r_{0}^{2}\sqrt{2(1-\sqrt{N_{0}})}r^{-2}+{\cal O}(r^{-6})$. Since the coupling (139) behaves as $\mu\to\mu_{\infty}={\rm constant}$ at large distances, [^5] Using the solution (124) for $h$ at large distances, we find $\mu\simeq M_{\rm Pl}^{-4n+2}n^{-2n}2^{-n}c_{2}\left(c_{2}^{2}/q_{E}^{2}\right)^{-n}\left(2n-1\right)^{4n-2}/\left(3-2n\right)^{2n-1}$, and hence $c_{2}$ needs to be non-negative. this translates to the $\phi$ dependence $\mu(\phi)=\mu(\phi_{\infty})$. By fixing $\phi_{\infty}$ to the cosmological value of $\phi$, one can uniquely determine the functional form of $\mu$, as both $\phi$ and $\mu$ are completely determined as functions of $r$.
V.2 Magnetic case
For $q_{M}\neq 0$ and $q_{E}=0$, we have $A_{0}^{\prime}=0$ and hence Eq. (10) is automatically satisfied. From Eq. (9), we have
$\phi^{\prime}=M_{\rm Pl}\sqrt{\frac{N^{\prime}}{rN}}\,,$ (142)
where we have chosen the branch $\phi^{\prime}>0$. The Lagrangian ${\cal L}$ obeys the following relation
$-\frac{M_{\rm Pl}^{2}hN^{\prime}}{2rN}+\mu(\phi)\left(-\frac{q_{M}^{2}}{2r^{4}}\right)^{n}=\frac{M_{\rm Pl}^{2}}{r^{2}}\left(rh^{\prime}+h-1\right)\,.$ (143)
If $n$ is an integer, the second term in Eq. (143) reduces to $[q_{M}^{2}/(2r^{4})]^{n}$ for even $n$ and $-[q_{M}^{2}/(2r^{4})]^{n}$ for odd $n$. If $n$ is not an integer, we should think of $F^{n}$ in the Lagrangian as $(F^{2})^{m}$, i.e., $n=2m$, so that Eq. (143) reduces to
$\mu=M_{\rm Pl}^{2}\left(\frac{4r^{8}}{q_{M}^{4}}\right)^{n/2}\left(\frac{rh^{\prime}-h-1}{r^{2}}+\frac{hN^{\prime}}{2rN}\right)\,.$ (144)
For both integer and non-integer values of $n$, the coupling $\mu(\phi)$ is known in terms of $h$, $N$, and their $r$ derivatives.
From Eq. (11), we can express $\mu_{,\phi}$, as
$\mu_{,\phi}=-\frac{M_{\rm Pl}[rhN^{\prime\prime}+(2rh^{\prime}+3h)N^{\prime}]}{2r^{3/2}\sqrt{NN^{\prime}}}\left(-\frac{q_{M}^{2}}{2r^{4}}\right)^{-n}\,.$ (145)
We take the $r$ derivative of Eq. (143) and exploit the relation $\mu^{\prime}=\mu_{,\phi}\phi^{\prime}$. Eliminating the term $\mu_{,\phi}$ on account of Eq. (105), it follows that
$2(r^{2}h^{\prime\prime}+4nrh^{\prime}+4nh-2h-4n+2)N^{2}-r^{2}hN^{\prime 2}$
$+(3rh^{\prime}+4nh)rNN^{\prime}+2rhN(rN^{\prime\prime}+N^{\prime})=0\,.$ (146)
This equation can be mapped to the differential Eq. (109) found for the pure electric case by the substitution $n\to n/(2n-1)$. We use the labels “e” and “m” for the electric and magnetic cases, respectively. If $n_{\rm e}>1/2$ for the absence of Laplacian instabilities, the mapping $n_{\rm m}=n_{\rm e}/(2n_{\rm e}-1)$ implies that $n_{\rm e}=n_{\rm m}/(2n_{\rm m}-1)$, and also $n_{\rm m}>1/2$. Therefore, we have the mirror magnetic solutions without Laplacian instabilities identified by using the above mapping. The same mapping brings Eq. (102) into Eq. (103) and vice versa.
The integrated solution to Eq. (146) can be obtained by replacing the power $n$ to $n/(2n-1)$ in Eq. (111), so that
$h(r)$ $=$ $\frac{2(2n-1)\int_{0}^{r}r_{1}^{4(n-1)}\sqrt{N(r_{1})}\bigl{(}\int_{0}^{r_{1}}\sqrt{N(r_{2})}{\rm d}r_{2}\bigr{)}{\rm d}r_{1}}{r^{2(2n-1)}N(r)}$ (147)
$+\frac{c_{1}}{r^{2(2n-1)}N(r)}$
$+\frac{c_{2}(2n-1)\int_{0}^{r}r_{1}^{4(n-1)}\sqrt{N(r_{1})}{\rm d}r_{1}}{r^{2(2n-1)}N(r)}\,.$
To avoid the divergence of $h(r)$ at $r=0$ for $n>1/2$, we require that $c_{1}=0=c_{2}$. Then, the resulting solution to $h(r)$ is given by
$h(r)=\frac{2(2n-1)\int_{0}^{r}r_{1}^{4(n-1)}\sqrt{N(r_{1})}\bigl{(}\int_{0}^{r_{1}}\sqrt{N(r_{2})}{\rm d}r_{2}\bigr{)}{\rm d}r_{1}}{r^{2(2n-1)}N(r)}\,.$ (148)
Since $h(r)>0$ at any $r\geq 0$ under the Laplacian stability condition $n>1/2$, there are no regular magnetic BHs with apparent horizons.
Figure 6: Metric component $h(r)$ versus $r/r_{0}$ for magnetic SSS objects in theories given by the Lagrangian (100). We choose $N(r)$ given by Eq. (120) with $N_{0}=0.1$. Each case corresponds to (a) $n=1/2$, (b) $n=4/5$, (c) $n=1$, (d) $n=2$, and (e) $n\gg 1$. The theoretical lines of $h(r)$ are positive at any distance $r$.
Figure 7: This plot depicts the allowed region of parameter space, indicated by the light-blue shaded area. The upper boundary corresponds to the value of $n_{\rm max}$, beyond which the no-ghost condition is violated for a given $N_{0}$ within the allowed range $10<N_{0}<$ 1. The lower boundary, at $n=3/4$, defines the limit below which the solutions no longer describe a compact object. Consequently, the permissible range for $n$ is $3/4<n\leq n_{\rm max}$.
To realize the behavior $\phi^{\prime}(r)\propto r$ for avoiding the cusp-like structure at $r=0$, we require that $N(r)$ is expanded as Eq. (115) around $r=0$. Then, the field derivative has the desired behavior (116). As a possible form of $N(r)$ having this property, we can choose the function (120). For this choice, $\phi^{\prime}(r)$ behaves as Eq. (131) at large distances. As we discussed for the electric case, $\phi^{\prime}(r)$ has a maximum at some intermediate distance, with the asymptotic behaviors $\phi^{\prime}(r)\propto r$ for $r\ll r_{0}$ and $\phi^{\prime}(r)\propto r^{-3}$ for $r\gg r_{0}$. We note that $A_{0}^{\prime}$ is vanishing at any distance $r$ for the magnetic configuration.
The metric component $h$ behaves as Eq. (118) around $r=0$, where $n$ should be replaced with $n/(2n-1)$. At large distances, we require that the ADM mass $M$ is converging, whose condition corresponds to $n<3/2$ for the electric configuration. Using the electric-magnetic duality, the condition for compactness of magnetic SSS objects is now given by $n>3/4$. For the choice (120), the large-distance solutions to $h(r)$ are given by the dual of Eq. (124), that is
$h(r)=1+\frac{c_{1}}{r}+\frac{c_{2}}{r^{4n-2}}=1+\frac{c_{1}}{r}+\mathcal{O}(r^{2-4n})\,.$ (149)
For the parameter range $n>3/4$, the term of order $r^{2-4n}$ is sub-dominant to $c_{1}/r$.
For $n\to 1/2$, we have the solution corresponding to the limit $n\to\infty$ in the electric configuration, i.e.,
$h_{n\to 1/2}(r)=\frac{N_{0}}{N(r)}\,.$ (150)
For $n\gg 1$, the solution is equivalent to that for $n=1/2$ in the electric case, i.e.,
$h_{n\gg 1}(r)=\frac{1}{r\sqrt{N(r)}}\int_{0}^{r}\sqrt{N(r_{1})}\,{\rm d}r_{1}\,.$ (151)
In Fig. 6, we show $h$ versus $r/r_{0}$ for $n=1/2,4/5,1,2$ and the limit $n\gg 1$, with the choice $N_{0}=0.1$ in Eq. (120). The theoretical curves are bounded from below by the line $h_{n\to 1/2}(r)$. We have $h(r)>0$ for all the powers $n$ in the range $n>1/2$, showing the absence of nonsingular magnetic BHs with apparent horizons. The compactness of magnetic SSS objects requires that $n>3/4$. In such cases, we numerically checked the asymptotic behaviors $h-1\propto r^{4}$ for $r\ll r_{0}$ and $h-1\propto r^{-1}$ for $r\gg r_{0}$. We note that the radial dependence of $\phi^{\prime}(r)$ is similar to that of the electric case shown in Fig. 3.
Using Eq. (143), the no-ghost condition ${\cal L}_{,F}=n\mu[-q_{M}^{2}/(2r^{4})]^{n-1}>0$ translates to
${\cal L}_{,F}=\frac{q_{E}^{2}(1-2n)N}{M_{\rm Pl}^{2}nr^{2}[rhN^{\prime}+(rh^{\prime}+h-1)N]}>0\,.$ (152)
In comparison to Eq. (138) for the electric case, there is an electric-magnetic duality in that Eq. (152) follows from the inverse of Eq. (138) by the mappings $n\to n/(2n-1)$ and $q_{E}\to q_{M}$. This means that we can derive the no-ghost condition for magnetic SSS objects by using the correspondence $n_{\rm e}=n_{\rm m}/(2n_{\rm m}-1)$. Then, in the magnetic configuration, the range (140) obtained for the electric case translates to the following region
$\frac{3}{4}<n\leq n_{\rm max}\equiv\frac{n_{{\rm e,min}}}{2n_{{\rm e,min}}-1}\,,$ (153)
where $n_{{\rm e,min}}$ is the minimum value of $n$ given in Eq. (140). In Fig. 7, we plot the allowed range (153) versus $N_{0}$ for the choice (120). We find that $n_{\rm max}$ is larger than 1, e.g. $n_{\rm max}=1.19$ for $N_{0}=0.1$. As $N_{0}$ increases in Fig. 7 from 0.01 to 1, $n_{\rm max}$ decreases from 1.253 to 1. In the parameter region shown as a shaded color, there are compact magnetic SSS objects with neither ghosts nor Laplacian instabilities.
From Eq. (143), the coupling $\mu$ can be written in terms of $h(r)$, $N(r)$, and their $r$ derivatives. Integrating Eq. (142), we can determine $\phi$ as a growing function of $r$. This allows us to express $\mu$ as a function of $\phi$. Since $F$ is negative, the no-ghost condition $\mathcal{L}_{,F}=n(\mu/F)F^{n}>0$ means that $\mu<0$ for even-integer $n$ and $\mu>0$ for odd-integer $n$. If $n$ is non-integer, we should consider $F^{n}$ as $(F^{2})^{m}$, so that $n\to 2m$. In this case, we have that $\mathcal{L}_{,F}=n(\mu/F)(F^{2})^{m}$, and the condition $\mathcal{L}_{,F}>0$ leads to $\mu/F>0$, i.e., $\mu<0$. Thus, depending on the values of $n$, the signs of $\mu$ consistent with no-ghost conditions are different.
VI Theories with Lagrangian $\mathcal{L}=X\kappa(F)$
Finally, we consider theories given by the Lagrangian
${\cal L}=X\kappa(F)\,,$ (154)
where $\kappa$ is a function of $F$. The Lagrangian (154) respects a shift symmetry for the scalar field, so that it is invariant under the shift $\partial_{\mu}\phi\to\partial_{\mu}\phi+c_{\mu}$. In this case, we have $\mathcal{L}_{,\phi}=0$ and hence Eq. (15) gives
$\phi^{\prime}=\frac{M_{\rm Pl}^{2}N^{\prime}hr}{q_{s}\sqrt{N}}\,,$ (155)
where $q_{s}$ is a constant. From Eq. (9), we obtain
$\phi^{\prime 2}=\frac{M_{\rm Pl}^{2}N^{\prime}}{\kappa rN}\,.$ (156)
Combining Eq. (156) with Eq. (155), we can express $\kappa$ in the form
$\kappa=\frac{q_{s}^{2}}{M_{\rm Pl}^{2}N^{\prime}h^{2}r^{3}}\,,$ (157)
which is a function of $r$.
VI.1 Electric case
Let us first study the case $q_{E}\neq 0$ and $q_{M}=0$. Since ${\cal L}_{,X}=\kappa$, ${\cal L}_{,XX}=0$, and ${\cal L}_{,FX}=\kappa_{,F}$, the squared propagation speeds (III.2.2) and (68) reduce, respectively, to
$c_{r,\delta\phi}^{2}$ $=$ $1-\frac{4F\kappa_{,F}^{2}}{\kappa(\kappa_{,F}+2F\kappa_{,FF})}\,,$ (158)
$c_{\Omega,V}^{2}$ $=$ $\frac{\kappa_{,F}}{\kappa_{,F}+2F\kappa_{,FF}}\,.$ (159)
We first derive the relation between $h(r)$ and $N(r)$ to find the existence of nonsingular BHs. Using Eq. (156), the Lagrangian (154) is expressed as
${\cal L}=-\frac{M_{\rm Pl}^{2}hN^{\prime}}{2rN}\,,$ (160)
on the SSS background. Equating Eq. (160) with Eq. (13) and using Eq. (21), it follows that
$2(r^{2}h^{\prime\prime}+2rh^{\prime})N^{2}-r^{2}hN^{\prime 2}+(3rh^{\prime}+2h)rNN^{\prime}$
$+2rhN(rN^{\prime\prime}+N^{\prime})=0\,.$ (161)
This differential equation is the same as Eq. (146) with $n=1/2$. Then, we have the following integrated solution
$h(r)=\frac{c_{1}}{N(r)}+\frac{c_{2}}{N(r)}\int_{0}^{r}\frac{\sqrt{N(r_{1})}}{r_{1}^{2}}{\rm d}r_{1}\,.$ (162)
Imposing regularities at $r=0$ in the forms (17) and (18), we can fix integration constants to be $c_{1}=N_{0}$ and $c_{2}=0$. Then, Eq. (162) reduces to
$h(r)=\frac{N_{0}}{N(r)}\,.$ (163)
Since $N(r)>0$ and $N_{0}>0$, $h(r)$ is positive at any distance $r$. This means that nonsingular electric BHs do not exist in theories given by the Lagrangian (154). In particular, by setting asymptotically-flat boundary conditions $h\to 1$ and $N\to 1$ for $r\to\infty$, we need to impose that $N_{0}=1$ and hence $h(r)=1/N(r)$. In this case, we have $h=N=1$ at both $r=0$ and $r\to\infty$, but $h(r)$ and $N(r)$ can differ from 1 at intermediate distances.
The above SSS configuration does not correspond to nonsingular BHs, but we study the linear stability of such SSS objects without horizons. From Eqs. (157) and (163), the coupling $\kappa$ can be expressed as
$\kappa=\frac{q_{s}^{2}N^{2}}{M_{\rm Pl}^{2}N_{0}^{2}r^{3}N^{\prime}}\,.$ (164)
On using this relation with $F=A_{0}^{\prime 2}/(2N)$, we can compute the quantities $\kappa_{,F}$ and $\kappa_{,FF}$ in Eqs. (158) and (159). Then, it follows that
$c_{r,\delta\phi}^{2}$ $=$ $\frac{NN^{\prime}}{2rN^{\prime 2}-N(rN^{\prime\prime}+2N^{\prime})}\,,$ (165)
$c_{\Omega,V}^{2}$ $=$ $\frac{rN_{0}N^{\prime}(rNN^{\prime\prime}-2rN^{\prime 2}+3NN^{\prime})}{2(2N_{0}N-rN_{0}{N^{\prime}}-2N^{2})(rNN^{\prime\prime}-2r{N^{\prime}}^{2}+2N{N^{\prime}})}\,.$ (166)
Using the expansion (18) of $N(r)$ around $r=0$, we obtain
$c_{r,\delta\phi}^{2}$ $=$ $-\frac{1}{3}+\frac{N_{3}}{6N_{2}}\,r+\mathcal{O}(r^{2})\,,$ (167)
$c_{\Omega,V}^{2}$ $=$ $-\frac{1}{3}-\frac{N_{3}}{24N_{2}}\,r+\mathcal{O}(r^{2})\,,$ (168)
whose leading-order terms are negative. Hence the background SSS solution is prone to Laplacian instability in both radial and angular directions. We stress that this instability arises for the coupling $\kappa$ with the radial dependence given by Eq. (164). In the vicinity of $r=0$, we have
$\kappa$ $=$ $\frac{q_{s}^{2}}{2M_{\rm Pl}^{2}N_{2}}r^{-4}+{\cal O}(r^{-3})\,,$ (169)
$F$ $=$ $\frac{2M_{\rm Pl}^{4}N_{2}^{2}}{q_{E}^{2}N_{0}^{2}}r^{4}+{\cal O}(r^{5})\,.$ (170)
Then, the leading-order terms of $\kappa$ and $F$ have the following relation
$\kappa=\frac{M_{\rm Pl}^{2}N_{2}q_{s}^{2}}{q_{E}^{2}N_{0}^{2}}F^{-1}\,,$ (171)
around $r=0$. The reason why the leading-order terms of $c_{r,\delta\phi}^{2}$ and $c_{\Omega,V}^{2}$ are negative is attributed to the property $\kappa\propto F^{-1}$. Indeed, we can obtain the value $-1/3$ by substituting the relation $\kappa\propto F^{-1}$ into Eqs. (158) and (159). Even though the dependence of $\kappa$ on $F$ is different for the distance $r$ away from the origin, the Laplacian instability around $r=0$ is sufficient to exclude the above horizonless solution as a stable SSS configuration.
VI.2 Magnetic case
For $q_{M}\neq 0$ and $q_{E}=0$, the squared radial and angular propagation speeds (85) and (86) reduce to
$c_{r,\delta\phi}^{2}$ $=$ $1\,,$ (172)
$c_{\Omega,\delta A}^{2}$ $=$ $\frac{\kappa_{,F}+2F\kappa_{,FF}}{\kappa_{,F}}\,.$ (173)
This means that the Laplacian instability of $\delta\phi$ along the radial direction is absent.
As in the electric case, the background Lagrangian is expressed as Eq. (160). Combing this with Eq. (13) and using the property $A_{0}^{\prime}=0$, we obtain
$2N\left(rh^{\prime}+h-1\right)=-rhN^{\prime}\,.$ (174)
This differential equation is equivalent to the limit $n\gg 1$ in Eq. (146). Then, the integrated solution to $h(r)$ that is regular at $r=0$ is given by
$h(r)=\frac{1}{r\sqrt{N(r)}}\int_{0}^{r}\sqrt{N(r_{1})}\,{\rm d}r_{1}\,,$ (175)
which is positive for $r\geq 0$. Then, the above solution does not correspond to the nonsingular BH with an apparent horizon.
Since the coupling $\kappa$ is given by Eq. (157), we can express $\kappa_{,F}$ and $\kappa_{,FF}$ in terms of $h(r)$, $N(r)$ and its $r$ derivatives. Then, Eq. (173) yields
$c_{\Omega,\delta A}^{2}$ $=-(2r^{2}h^{2}NN^{\prime\prime 2}-r^{2}h^{2}NN^{\prime}N^{\prime\prime\prime}-r^{2}h^{2}N^{\prime 2}N^{\prime\prime}-rh^{2}NN^{\prime}N^{\prime\prime}+rh^{2}N^{\prime 3}-h^{2}NN^{\prime 2}+4rhNN^{\prime}N^{\prime\prime}-5rhN^{\prime 3}$
$~{}~{}~{}-2hNN^{\prime 2}+6NN^{\prime 2})/[2h(rhN^{\prime 2}-rhNN^{\prime\prime}-hNN^{\prime}-2NN^{\prime})N^{\prime}]\,.$ (176)
Using the expansions (17) and (18) of $h(r)$ and $N(r)$ around $r=0$, it follows that
$c_{\Omega,\delta A}^{2}=1+\frac{9N_{3}}{16N_{2}}\,r+{\cal O}(r^{2})\,.$ (177)
Since the leading-order contribution to $c_{\Omega,\delta A}^{2}$ is positive, the angular Laplacian instability of $\delta A$ is absent around $r=0$. On the other hand, at large distances, we exploit the expanded solutions of $N(r)$ and $h(r)$ in the forms
$N(r)=\sum_{n=0}\frac{\tilde{N}_{n}}{r^{n}}\,,\qquad h(r)=1+\sum_{n=1}\frac{\tilde{h}_{n}}{r^{n}}\,,$ (178)
where $\tilde{N}_{n}$ and $\tilde{h}_{n}$ are constants. Solving Eq. (174) order by order, we have $\tilde{N}_{1}=0$ and $\tilde{h}_{2}=-\tilde{N}_{2}/\tilde{N}_{0}$ at lowest order. This leaves $\tilde{h}_{1}$ as a free parameter, related to the ADM mass of the SSS object.
Then, the large-distance behavior of $c_{\Omega,\delta A}^{2}$ is given by
$c_{\Omega,\delta A}^{2}=-\frac{1}{2}+\frac{8\tilde{N}_{0}\tilde{N}_{2}\tilde{N}_{4}-12\tilde{N}_{0}\tilde{N}_{2}^{2}\tilde{h}_{1}^{2}-12\tilde{N}_{0}\tilde{N}_{2}\tilde{N}_{3}\tilde{h}_{1}-9\tilde{N}_{0}\tilde{N}_{3}^{2}-8\tilde{N}_{2}^{3}}{2\tilde{N}_{0}\tilde{N}_{2}\,(4\tilde{N}_{2}\tilde{h}_{1}+3\tilde{N}_{3})}\,\frac{1}{r}+{\cal O}(r^{-2})\,,$ (179)
whose leading-order term is negative. Hence the nonsingular SSS object without the horizon is excluded by angular Laplacian instability at large distances. We note that, even without imposing the condition $\tilde{N}_{0}=1$ for asymptotic flatness, the leading-order contribution to $c_{\Omega,\delta A}^{2}$ is $-1/2$. The above results show that, as $r$ increases from 0, $c_{\Omega,\delta A}^{2}$ enters the regime $c_{\Omega,\delta A}^{2}<0$ to approach the asymptotic value $-1/2$. Indeed, we have numerically confirmed this property [^6] This behavior of $c_{\Omega,\delta A}^{2}$ is similar to what happens for hairy BHs present in cubic vector Galileon theories [64]. by choosing $N(r)$ of the form (120).
VII Conclusions
In this paper, we extended our previous analysis of the nonsingular BHs in NED [42] to more general theories characterized by the matter Lagrangian ${\cal L}(F,\phi,X)$ with an Einstein-Hilbert term. In NED with the Lagrangian ${\cal L}(F)$, the charged SSS objects that are nonsingular at their centers exhibit angular Laplacian instabilities arising from vector-field perturbations. The primary objective of our study was to investigate whether this property persists when a scalar field $\phi$ is introduced into the theory.
In Sec. II, we first showed that the background solutions with mixed electric and magnetic charges do not exist. Then, as in the case of NED, we could separate the analysis into either electrically or magnetically charged objects. For the electric configuration, both $A_{0}^{\prime}$ and $\phi^{\prime}$ are generally nonvanishing, while $A_{0}^{\prime}=0$ for the magnetic configuration.
In Sec. III, we derived the second-order action of perturbations by taking into account both electric and magnetic charges. For the electric configuration, the action consists of dynamical perturbations in the odd-parity sector A and the even-parity sector B. In this case, we showed that there are neither ghosts nor Laplacian instabilities under the conditions ${\cal L}_{,X}>0$, ${\cal L}_{,F}>0$, $c_{r,\delta\phi}^{2}>0$, and $c_{\Omega,\delta A}^{2}>0$, where $c_{r,\delta\phi}^{2}$ and $c_{\Omega,V}^{2}$ are given, respectively by Eqs. (III.2.2) and (68). For the magnetic case, the action can be decomposed into two sectors C and D, both of which contain the contributions of odd- and even-parity dynamical perturbations. In this case, we obtained the linear stability conditions ${\cal L}_{,X}>0$, ${\cal L}_{,F}>0$, $c_{r,\delta\phi}^{2}>0$, and $c_{\Omega,\delta A}^{2}>0$, where $c_{r,\delta\phi}^{2}$ and $c_{\Omega,\delta A}^{2}$ are given, respectively by Eqs. (85) and (86).
In Sec. IV, we applied the linear stability conditions to NED with a k-essence scalar field. For nonsingular electric SSS objects, we showed that the even-parity vector-field perturbation $V$ is subject to angular Laplacian instability around $r=0$ due to the negative leading term of $c_{\Omega,V}^{2}$. For the nonsingular magnetic configuration, the same angular Laplacian instability arises for the odd-parity vector-field perturbation $\delta A$. Since the fields $V$ and $\delta A$ are coupled to the even-parity gravitational perturbation $\psi$, the background metric regular at $r=0$ cannot remain in a steady state. This means that, in theories given by the Lagrangian ~$\tilde{\cal L}(F)+K(\phi,X)$, the nonsingular SSS objects (including regular BHs) do not exist as stable configurations.
In Sec. V, we analyzed regular solutions for the Lagrangian of the form ${\cal L}=X+\mu(\phi)F^{n}$. For both electric and magnetic SSS objects, we showed that the absence of angular Laplacian instabilities imposes the condition $n>1/2$. Under this inequality, the metric components are positive at any distance $r$, so that there are no regular BHs with apparent horizons.
In Sec. V, we also studied the linear stability of nonsingular horizonless SSS objects further. For a given function of $N(r)$, the profile of $\phi(r)$ is identical for both the electric and magnetic configurations. To avoid the formation of cusp-like structures, we impose the regular behavior $\phi^{\prime}(r)\propto r$ around $r=0$. One of the examples for $N(r)$ consistent with this boundary condition and asymptotic flatness is given by Eq. (120).
For the electric horizonless SSS configuration studied in Sec. V, the solutions can be described by regular compact objects if $n<3/2$. In this case, the absence of ghosts requires that $n_{\rm min}\leq n$, where $n_{\rm min}$ is the minimum value of $n$ larger than $1/2$. As shown in Fig. 5, $n_{\rm min}$ mildly depends on $N_{0}$, where $N_{0}$ is a constant appearing in Eq. (120). Therefore, the allowed parameter space of electric regular compact objects is given by $n_{\rm min}\leq n<3/2$. Since there is a duality relation $n_{\rm e}=n_{\rm m}/(2n_{\rm m}-1)$ between electric and magnetic cases, the existence of linearly stable magnetic compact objects requires the condition $3/4<n\leq n_{\rm max}=n_{\rm e,min}/(2n_{\rm e,min}-1)$.
Finally, we investigated Lagrangians of the form ${\cal L}=X\kappa(F)$ in Sec. VI. Similar to the previous case, these theories do not support the SSS configurations with apparent horizons, precluding the existence of nonsingular BHs. Unlike theories discussed in Sec. V, all regular SSS solutions obtained in this framework (including those without horizons) suffer from angular Laplacian instabilities. Specifically, for electric solutions, such instabilities arise around the regular center. In contrast, for magnetic solutions, they manifest at large distances away from the origin.
To summarize, we have not found any linearly stable nonsingular BHs for theories studied in this paper. To reach this conclusion, we have made three assumptions: (i) the choice of three subsets of ${\cal L}(F,\phi,X)$ theories, (ii) the spherically symmetric and static regular background with asymptotic flatness, and (iii) the existence of the BH apparent horizon. To circumvent the no-go result, one could either modify the theories—by extending the analysis to other subclasses of ${\cal L}(F,\phi,X)$ theories and more general theories—or alter the spacetime background. For instance, rotating regular BH solutions, if they exist, might offer a stable alternative.
Our results show that the construction of regular BHs in the framework of classical field theories is highly complex and challenging. In other words, Penrose’s singularity theorem holds even in scenarios where some of its original assumptions are violated. If this property generally persists for local classical four-dimensional actions, we may need to resort to non-local theories of gravity or higher-dimensional theories (see, e.g., [65, 66, 67, 68, 69, 70]). On the other hand, our analysis in Sec. V has uncovered linearly stable regular SSS solutions without horizons that, in principle, could correspond to physically realizable configurations in nature. The exploration of physical and geometric properties of such stable compact objects without horizons, including the comparisons with boson and Proca stars [71, 72], remains an intriguing avenue for future research.
Acknowledgements
We thank Pablo Cano, Vitor Cardoso, Hiroki Takeda, and Kent Yagi for useful discussions. The Japan Society for the Promotion of Science supported ADF’s work through Grants-in-Aid for Scientific Research No. 20K03969. ST was supported by the Grant-in-Aid for Scientific Research Fund of the JSPS No. 22K03642 and Waseda University Special Research Project No. 2024C-474.
Appendix A: Coefficients of the second-order action
The explicit forms of coefficients appearing in the Lagrangians (33) and (34) are given by
$a_{0}=\frac{r^{2}A_{0}^{\prime 2}}{8N^{3/2}}(N{\cal L}_{,F}+A_{0}^{\prime 2}{\cal L}_{,FF}),\qquad a_{1}=-\frac{M_{\rm Pl}^{2}rh\sqrt{N}}{2},\qquad a_{2}=\frac{M_{\rm Pl}^{2}h\sqrt{N}}{2},$
$a_{3}=\frac{Nr^{2}A_{0}^{\prime 2}({\cal L}_{,F}+h\phi^{\prime 2}{\cal L}_{,XF})-r^{2}A_{0}^{\prime 4}{\cal L}_{,FF}-N^{2}[2M_{\rm Pl}^{2}+r^{2}(2{\cal L}+h\phi^{\prime 2}{\cal L}_{,X})]}{4N^{3/2}}\qquad a_{4}=-\frac{M_{\rm Pl}^{2}\sqrt{N}}{4},$
$a_{5}=\frac{M_{\rm Pl}^{2}N(h+1)+r^{2}(N{\cal L}-A_{0}^{\prime 2}{\cal L}_{,F})}{4\sqrt{N}r},\qquad a_{6}=\frac{q_{M}(N{\cal L}_{,F}-A_{0}^{\prime 2}{\cal L}_{,FF})}{2\sqrt{N}\,r^{2}},\qquad a_{7}=\frac{hr^{2}\phi^{\prime}(N{\cal L}_{,X}-A_{0}^{\prime 2}{\cal L}_{,XF})}{2\sqrt{N}},$
$a_{8}=\frac{r^{2}(A_{0}^{\prime 2}{\cal L}_{,\phi F}-N{\cal L}_{,\phi})}{2\sqrt{N}},\qquad b_{1}=\frac{M_{\rm Pl}^{2}}{4\sqrt{N}},\qquad b_{2}=\frac{M_{\rm Pl}^{2}r}{\sqrt{N}},\qquad b_{3}=-2b_{1}\,,\qquad b_{4}=-\frac{r^{2}\phi^{\prime}{\cal L}_{,X}}{\sqrt{N}}\,,$
$c_{0}=-\frac{a_{3}}{2}+\frac{hr^{2}\phi^{\prime 2}(h\phi^{\prime 2}N{\cal L}_{,XX}-A_{0}^{\prime 2}{\cal L}_{,XF})}{8\sqrt{N}},\qquad c_{1}=-a_{5}-\frac{hr\phi^{\prime 2}\sqrt{N}{\cal L}_{,X}}{4},\qquad c_{2}=-a_{6}-\frac{q_{M}h\phi^{\prime 2}\sqrt{N}{\cal L}_{,XF}}{2r^{2}},$
$c_{3}=\frac{hr^{2}\phi^{\prime}[N({\cal L}_{,X}-h\phi^{\prime 2}{\cal L}_{,XX})+A_{0}^{\prime 2}{\cal L}_{,XF}]}{2\sqrt{N}},\qquad c_{4}=\frac{r^{2}[N({\cal L}_{,\phi}+h\phi^{\prime 2}{\cal L}_{,\phi X})-A_{0}^{\prime 2}{\cal L}_{,\phi F}]}{2\sqrt{N}},\qquad$
$d_{0}=b_{1},\qquad d_{1}=\frac{h\sqrt{N}(M_{\rm Pl}^{2}r^{2}-q_{M}^{2}{\cal L}_{,F})}{2r^{4}},\qquad d_{2}=-\frac{A_{0}^{\prime}{\cal L}_{,F}}{\sqrt{N}},\qquad d_{3}=-\frac{q_{M}h\sqrt{N}{\cal L}_{,F}}{r^{2}},\qquad d_{4}=h\phi^{\prime}\sqrt{N}{\cal L}_{,X}\,,$
$s_{1}=\frac{r^{2}(N{\cal L}_{,F}+A_{0}^{\prime 2}{\cal L}_{,FF})}{2N^{3/2}},\qquad s_{2}=A_{0}^{\prime}s_{1},\qquad s_{3}=-s_{2}+\frac{hr^{2}\phi^{\prime 2}A_{0}^{\prime}{\cal L}_{,XF}}{2\sqrt{N}},\qquad s_{4}=-\frac{q_{M}A_{0}^{\prime}{\cal L}_{,FF}}{\sqrt{N}\,r^{2}},$
$s_{5}=-\frac{hr^{2}\phi^{\prime}A_{0}^{\prime}{\cal L}_{,XF}}{\sqrt{N}},\qquad s_{6}=\frac{r^{2}A_{0}^{\prime}{\cal L}_{,\phi F}}{\sqrt{N}},\qquad s_{7}=\frac{{\cal L}_{,F}}{2h\sqrt{N}},\qquad s_{8}=-\frac{h\sqrt{N}{\cal L}_{,F}}{2}\,,$
$u_{1}=\frac{r^{2}{\cal L}_{,X}}{2h\sqrt{N}},\qquad u_{2}=\frac{r^{2}h\sqrt{N}(h\phi^{\prime 2}{\cal L}_{,XX}-{\cal L}_{,X})}{2},\qquad u_{3}=-\frac{\sqrt{N}{\cal L}_{,X}}{2},\qquad\tilde{u}_{3}=\frac{1}{2}\frac{\partial{\cal E}_{\phi}}{\partial\phi},$
$u_{4}=\frac{Lq_{M}h\phi^{\prime}\sqrt{N}{\cal L}_{,XF}}{r^{2}},\qquad u_{5}=-\frac{Lq_{M}\sqrt{N}{\cal L}_{,\phi F}}{r^{2}}\,,$ (A.1)
$p_{1}=\frac{M_{\rm Pl}^{2}}{4\sqrt{N}\,r^{2}},\qquad p_{2}=\frac{d_{2}}{r},\qquad p_{3}=s_{7},\qquad p_{4}=s_{8},\qquad p_{5}=\frac{\sqrt{N}(q_{M}^{2}{\cal L}_{,FF}-r^{4}{\cal L}_{,F})}{2r^{6}},\qquad p_{6}=-\frac{M_{\rm Pl}^{2}h\sqrt{N}}{4r^{2}},$
$p_{7}=d_{1},\qquad p_{8}=\frac{M_{\rm Pl}^{2}}{4r^{2}h\sqrt{N}},\qquad p_{9}=-\frac{d_{1}}{h^{2}N},\qquad p_{10}=-\frac{d_{3}}{h^{2}N},\qquad p_{11}=-A_{0}^{\prime}hp_{10},\qquad p_{12}=-h^{2}Np_{10}\,,$ (A.2)
where ${\cal E}_{\phi}$ is defined in Eq. (11).